### v2

# A Structural Characterization of the Hit Image in the Motivic Steenrod Algebra

Abstract.
The motivic hit problem seeks a minimal generating set for the motivic cohomology of classifying spaces as a module over the motivic Steenrod algebra. In this paper, we provide a complete structural classification of the hit image within the top layer of Kameko’s decomposition, revealing a sharp dichotomy governed by parity. Specifically, for degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, let $V$ denote the span of the monotone translates of Kameko’s monomial $z_{k}$. We prove that the intersection of $V$ with the hit subspace is exactly the even-parity hyperplane. Consequently, the quotient $V/(\text{hits})$ is one-dimensional and is generated by any odd-parity linear combination of these translates. This establishes the ”odd-parity gap” as the precise obstruction to the hit property in this layer. As an arithmetic consequence of this structural result, we identify a new infinite numerical family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$) where $\beta(d)>n$. This yields a new class of counterexamples to the motivic Peterson-type conjecture, distinct from Kameko’s original family. Furthermore, we show that these structural results and counterexamples persist over any algebraically closed field of characteristic 0.
Key words and phrases:
Motivic cohomology, motivic Steenrod algebra, hit problem, Peterson conjecture, parity functional
2020 Mathematics Subject Classification:
Primary 14F42, 55S10; Secondary 55S05, 55T15
1. Introduction
The classical hit problem asks for a minimal set of generators of the polynomial algebra
$P_{n}=\mathbb{F}_{2}[x_{1},\dots,x_{n}]$
as a module over the mod $2$ Steenrod algebra $A$; equivalently, it asks for a basis of the quotient
$QP_{n}^{d}:=P_{n}^{d}/(A^{+}(P_{n})\cap P_{n}^{d}),$
where $A^{+}\subset A$ denotes the ideal of positive-degree elements and $P_{n}^{d}$ is the homogeneous degree–$d$ part of $P_{n}$. This problem goes back to work of Peterson and Wood; see for instance the survey in Walker–Wood [9].
A key numerical invariant is the function
$\beta(d)=\min\Bigl\{s\in\mathbb{N}\,\Big|\,d=(2^{i_{1}}-1)+\cdots+(2^{i_{s}}-1),\ i_{j}\in\mathbb{N}\Bigr\},$
which already appears in Wood’s proof of the Peterson conjecture. Wood proved the following fundamental theorem [10].
Theorem 1.1 (Wood).
If $\beta(d)>n$, then $QP_{n}^{d}=0$.
Theorem 1.1 says that $\beta(d)>n$ is a sufficient condition for triviality of the hit quotient. Moreover, the condition is also necessary in the sense that the monomial $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ never lies in $A^{+}(P_{n})$. The interplay between the hit problem and the cohomology of $A$ via Singer’s transfer [7] connects this story to the classical Adams spectral sequence and the stable homotopy groups of spheres; see, for example, Ault [1], Pengelley–Williams [4], the present author [5, 6].
In motivic homotopy theory, Voevodsky introduced a motivic Steenrod algebra $A^{*,*}$ and reduced power operations in motivic cohomology over a base field of characteristic different from $2$ [8]. Prompted by the work of Dugger–Isaksen on the motivic Adams spectral sequence [2], it is natural to seek motivic analogues of the hit problem and of Peterson’s conjecture. Kameko initiated this study in [3], where he worked over the base field $\mathbb{C}$ and defined, for each $n\geq 1$, the bigraded ring
$M_{n}(\mathbb{C}):=H^{*,*}(BV_{n}/\mathbb{C};\mathbb{F}_{2})$
together with its $A^{*,*}$–module structure. He then formulated the motivic hit problem for the quotients
$QM_{n}^{d,*}(\mathbb{C}):=M_{n}(\mathbb{C})_{d,*}/(A^{>0,*}\cdot M_{n}(\mathbb{C})\cap M_{n}(\mathbb{C})_{d,*}),$
where $A^{>0,*}$ is the subspace of $A^{*,*}$ consisting of elements of positive total degree (topological plus weight). In direct analogy with Theorem 1.1, he proposed the following motivic version of Peterson’s conjecture.
Definition 1.2 (Motivic Peterson-type conjecture [3, Conj. 1.2]).
Assume the base field is $\mathbb{C}$. The motivic Peterson conjecture asserts that
$\beta(d)>n\quad\Longrightarrow\quad QM_{n}^{d,*}(\mathbb{C})=0$
for all $n,d$.
Kameko showed that this conjecture fails in general. For each $1\leq k<n$ he defined an explicit monomial $z_{k}\in M_{n}(\mathbb{C})$ and proved that $z_{k}$ is not in the image of the motivic Steenrod algebra. Moreover, when k=n3k=n-3 and $\alpha(n-2)\geq 3$ (where $\alpha(t)$ denotes the number of $1$’s in the binary expansion of $t$), he showed that $\beta(d)>n$ for the degree
$d=(n-1)\bigl(2^{k+1}-2\bigr)+k,$
so that Conjecture 1.2 fails in those degrees. More precisely, we recall:
Theorem 1.3 (Kameko [3, Thm. 1.3, Props. 1.4–1.5]).
Assume the base field is $\mathbb{C}$. For each $n\geq 1$ and $1\leq k<n$ there is a monomial $z_{k}\in M_{n}(\mathbb{C})$ of degree $d=(n-1)(2^{k+1}-2)+k$ such that $z_{k}\notin A^{>0,*}\cdot M_{n}(\mathbb{C})$. If in addition $\alpha(n-2)\geq 3$ and k=n3k=n-3, then $\beta(d)>n$, so that $QM_{n}^{d,*}(\mathbb{C})\neq 0$ despite $\beta(d)>n$.
Kameko’s proof is quite delicate. He passes to the quotient $N_{n}(\mathbb{C})=M_{n}(\mathbb{C})/(\tau)$ and constructs a family of monomials in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ that give a basis for the top layer of the motivic hit quotient. He then shows that only certain linear combinations can appear in the image of the Bockstein $Q_{0}$, and that $z_{k}$ itself is not among them.
The main objective of this paper is to uncover the sharp structural dichotomy governing the top layer of the motivic hit problem. While previous work focused on the existence of counterexamples, we provide a complete linear-algebraic characterization of the hit image in terms of parity.
  • (1)
    Structural Classification: We isolate a parity functional on the top layer of Kameko’s decomposition and prove that the hit image is exactly the even-parity hyperplane. This implies that the obstruction to the hit problem in this layer is one-dimensional and is completely determined by parity.
  • (2)
    Arithmetic Consequences: We apply this structural insight to the arithmetic of degrees $d=(n-1)(2^{k+1}-2)+k$. Using the equivalence between $\beta(d)>n$ and $\alpha(d+n)>n$ [3, Prop. 2.1], we identify a new infinite family of pairs $(n,k)$ with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) where the hit quotient is non-zero despite $\beta(d)>n$.
  • (3)
    Generalization: We demonstrate that these results are not specific to $\mathbb{C}$ but persist over any algebraically closed field $K$ of characteristic 0, relying on the stability of the motivic Steenrod algebra under base change.
Our main structural result is the following, which establishes that parity is the precise obstruction to being hit in the summand $V$.
Theorem 1.4 (Structure of the Top Hit Layer).
For any integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ be the degree of Kameko’s monomial $z_{k}$. Let $M_{1}$ be the set of all translates $\sigma(z_{k})$ under monotone permutations $\sigma\in\mathrm{Mono}(k)$, and let $V$ be their span in the motivic hit quotient $QN_{n}^{d,*}$.
Let $\theta:N_{n}^{d,*}\to V$ be the map extracting the $M_{1}$–component (defined precisely in Section 3). Then there is a linear functional
$\varepsilon:V\to\mathbb{F}_{2}$
such that:
  1. (1)
    Exact Image Characterization: One has
    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon)\subset V.$
  2. (2)
    Non-hit Criterion: Consequently, any class whose $M_{1}$–component has odd parity is non-hit. The quotient space V/(hit image)V/($\text${hit image}) is exactly one-dimensional.
Remark 1.5 (Minimality of the Obstruction).
Theorem 1.4 shows that the description of the hit image as the even-parity hyperplane is sharp. The “odd-parity gap” is the smallest possible non-trivial quotient, implying that no further reduction by the Steenrod algebra is possible in this layer.
On the arithmetic side, we prove:
Theorem 1.6 (New numerical family with $\beta(d)>n$).
Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let
$d=(n-1)\bigl(2^{k+1}-2\bigr)+k$
be the degree of $z_{k}$. Then $\beta(d)>n$.
Combining Theorems 1.3, 1.4 and 1.6 with the base-change discussion in Section 5, we obtain the following.
Theorem 1.7.
Let $K$ be an algebraically closed field of characteristic 0. For each integer $r\geq 5$, set $n=2^{r}+1$ and k=n4k=n-4, and let $d=(n-1)(2^{k+1}-2)+k$. Then:
  1. (1)
    $\beta(d)>n$.
  2. (2)
    The motivic hit quotient $QM_{n}^{d,*}(K)$ is non-zero.
  3. (3)
    More precisely, any odd-parity linear combination of the monomials $\sigma(z_{k})\in M_{n}^{d,*}(K)$ with $\sigma\in\mathrm{Mono}(k)$ represents a non-zero class in $QM_{n}^{d,*}(K)$.
Thus, combining the parity description of the top-layer hit image with the numerical condition $\beta(d)>n$, we obtain an infinite family of motivic Peterson-type counterexamples over any algebraically closed field of characteristic 0, distinct from Kameko’s k=n3k=n-3 family [3].
The paper is organized as follows. In Section 2 we review the classical hit problem, the motivic Steenrod algebra, and Kameko’s description of the motivic cohomology of $BV_{n}$ and the motivic hit problem. Section 3 introduces the parity functional on the top layer of the motivic hit quotient and proves Theorem 1.4. Section 4 is devoted to the arithmetic of the degrees $d=(n-1)(2^{k+1}-2)+k$, including Kameko’s original estimates and the new family in Theorem 1.6. Finally, Section 5 discusses the extension to arbitrary algebraically closed base fields of characteristic 0 and proves Theorem 1.7.
2. Preliminaries
In this section we collect definitions and known results needed later. We mostly follow the notation of Kameko [3].
2.1. The classical hit problem and the Peterson conjecture
Let $P_{n}=\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ be the graded polynomial algebra with $|x_{i}|=1$. Let $A$ be the mod $2$ Steenrod algebra, and let $A^{+}\subset A$ be the ideal of elements of positive degree. Following the standard convention, we define
$A^{+}(P_{n})=\{a\cdot v\mid a\in A^{+},\ v\in P_{n}\}.$
The hit problem asks for a basis of the quotient
$QP_{n}^{d}=P_{n}^{d}/(A^{+}(P_{n})\cap P_{n}^{d})$
in each degree $d$. The numerical invariant $\beta(d)$ introduced in the introduction is closely related to the function $\alpha(t)$ counting $1$’s in the binary expansion of $t$.
Proposition 2.1 (Kameko [3, Prop. 2.1]).
For positive integers $d,n$ one has $\beta(d)>n$ if and only if $\alpha(d+n)>n$.
Wood proved that if $\beta(d)>n$ then $QP_{n}^{d}=0$; see Theorem 1.1. The condition is also necessary in the sense that monomials of the form $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ are never hit.
2.2. Motivic cohomology of $BV_{n}$ and the motivic Steenrod algebra
Let $k$ be a field of characteristic not equal to $2$. Motivic cohomology with $\mathbb{F}_{2}$–coefficients is a bigraded ring $H^{p,q}(-;\mathbb{F}_{2})$ defined on smooth schemes over $k$; we refer to Voevodsky [8] for the construction. The motivic Steenrod algebra $A^{*,*}$ is a bigraded Hopf algebra acting on motivic cohomology, generated by a Bockstein $Q_{0}$ of bidegree $(1,0)$ and reduced power operations $\mathcal{P}^{a}$ of bidegree $(2a,a)$ for $a\geq 1$, subject to Adem and instability relations. Over fields of characteristic zero, the algebra $A^{*,*}$ and its action on $H^{*,*}(-;\mathbb{F}_{2})$ are described in detail in [2, 8].
Let $V_{n}=(\mathbb{Z}/2)^{n}$ and write $BV_{n}$ for the corresponding classifying space. Over $\mathbb{C}$, Kameko shows that there is an isomorphism of bigraded rings
(2.1) $M_{n}(\mathbb{C}):=H^{*,*}(BV_{n}/\mathbb{C};\mathbb{F}_{2})\cong\mathbb{F}_{2}[\tau,x_{1},\dots,x_{n},y_{1},\dots,y_{n}]/(x_{1}^{2}+\tau y_{1},\dots,x_{n}^{2}+\tau y_{n}),$
where $|\tau|=(0,1)$, $|x_{i}|=(1,1)$, and $|y_{i}|=(2,1)$ [3, Sec. 1]. The mod $2$ motivic Steenrod algebra $A^{*,*}$ is generated by $Q_{0}$ and the $\mathcal{P}^{a}$ with actions [3, Sec. 1]:
$Q_{0}(\tau)$ $=0,$ $\mathcal{P}^{a}(\tau)$ $=0$ $(a\geq 1),$
$Q_{0}(x_{i})$ $=y_{i},$ $\mathcal{P}^{a}(x_{i})$ $=0$ $(a\geq 1),$
$Q_{0}(y_{i})$ $=0,$ $\mathcal{P}^{1}(y_{i})$ $=y_{i}^{2},\quad\mathcal{P}^{a}(y_{i})=0$ $(a\geq 2),$
together with the Cartan formulas.
It is convenient to pass to the quotient
$N_{n}(\mathbb{C}):=M_{n}(\mathbb{C})/(\tau)\cong\Lambda(x_{1},\dots,x_{n})\otimes\mathbb{F}_{2}[y_{1},\dots,y_{n}],$
and to the quotient $A^{\sharp}:=A^{*,*}/(\tau)$, which is again generated by $Q_{0}$ and the reduced powers $\mathcal{P}^{a}$. We denote by $A^{\sharp}_{+}$ the subspace of elements of positive total degree (topological plus weight). We write $A^{\sharp}_{+}(N_{n})=\{a\cdot u\mid a\in A^{\sharp}_{+},\ u\in N_{n}\}$.
2.3. The motivic hit problem and the monomials $z_{k}$
Following Kameko, we define the motivic hit quotients as follows.
Definition 2.2 (Motivic hit quotients [3, Sec. 1]).
For each topological degree $d$ let $M_{n}(\mathbb{C})_{d,*}=\bigoplus_{q}H^{d,q}(BV_{n}/\mathbb{C};\mathbb{F}_{2})$. Set
$QM_{n}^{d,*}(\mathbb{C}):=M_{n}(\mathbb{C})_{d,*}/(A^{>0,*}\cdot M_{n}(\mathbb{C})\cap M_{n}(\mathbb{C})_{d,*}),$
where $A^{>0,*}\subset A^{*,*}$ is the subspace of elements of positive total degree. The problem of describing a basis for the $QM_{n}^{d,*}(\mathbb{C})$ is the motivic hit problem.
The quotient $N_{n}(\mathbb{C})$ is an $A^{\sharp}$–module, and there is an isomorphism
$QN_{n}^{d,*}(\mathbb{C}):=N_{n}(\mathbb{C})_{d,*}/\bigl(A^{\sharp}_{+}(N_{n}(\mathbb{C}))\cap N_{n}(\mathbb{C})_{d,*}\bigr)\cong QM_{n}^{d,*}(\mathbb{C})/\tau\,QM_{n}^{d,*}(\mathbb{C}),$
so we can work interchangeably with $QM_{n}^{d,*}(\mathbb{C})$ and $QN_{n}^{d,*}(\mathbb{C})$.
Kameko defines the function $\beta(d)$ exactly as in the classical case. As in (2.1), there are canonical monomials
$x_{1}\cdots x_{n}\,y_{1}^{2^{i_{1}-1}-1}\cdots y_{n}^{2^{i_{n}-1}-1}\in M_{n}(\mathbb{C})_{d,*}$
which are never in $A^{>0,*}\cdot M_{n}(\mathbb{C})$ when $\beta(d)\leq n$. This motivates the motivic Peterson-type conjecture (Conjecture 1.2).
For each pair of integers $n,k$ with $1\leq k<n$, Kameko defines a special monomial $z_{k}\in M_{n}(\mathbb{C})$ by
(2.2) $z_{k}=x_{1}\cdots x_{k}\left(\prod_{j=1}^{k}y_{j}^{2^{k}-2^{k-j}-1}\right)\left(\prod_{j=k+1}^{n}y_{j}^{2^{k}-1}\right),$
which has topological degree
(2.3) $d=(n-1)\bigl(2^{k+1}-2\bigr)+k.$
He proves:
Proposition 2.3 (Kameko [3, Prop. 1.4]).
For any integers $n\geq 1$ and $1\leq k<n$, the monomial $z_{k}$ defined in (2.2) is not in $A^{>0,*}\cdot M_{n}(\mathbb{C})$.
The key idea is to analyze the image of $Q_{0}$ in the quotient $N_{n}(\mathbb{C})/H_{n}$ for a certain subspace $H_{n}\subset N_{n}(\mathbb{C})$, and to describe the top layer of the hit quotient in terms of a pair of sets of monomials $M_{0}$ and $M_{1}$ which we review next.
2.4. Weight vectors and the subspaces $F_{n}$, $G_{n}$, and $H_{n}$
We briefly recall Kameko’s weight notation. Let $z=x_{1}^{\varepsilon_{1}}\cdots x_{n}^{\varepsilon_{n}}y_{1}^{e_{1}}\cdots y_{n}^{e_{n}}$ be a monomial in $N_{n}(\mathbb{C})$. Write
$\varepsilon_{i}+2e_{i}=\sum_{j\geq 0}\alpha_{ij}(z)2^{j},\qquad\alpha_{ij}(z)\in\{0,1\},$
and set
$\alpha_{i}(z)=\sum_{j\geq 0}\alpha_{ij}(z),\qquad\omega_{j}(z)=\sum_{i=1}^{n}\alpha_{ij}(z).$
Thus $\alpha_{i}(z)$ measures the number of non-zero bits in the binary expansion of $\varepsilon_{i}+2e_{i}$ and $\omega_{j}(z)$ measures how many of the $\varepsilon_{i}+2e_{i}$ have a $1$ in the $2^{j}$ place. We collect these into finite sequences
$\alpha(z)=(\alpha_{1}(z),\dots,\alpha_{n}(z)),\qquad\omega(z)=(\omega_{0}(z),\omega_{1}(z),\dots).$
Kameko defines a subspace $G_{n}\subset Y_{n}=\mathbb{F}_{2}[y_{1},\dots,y_{n}]$ generated by:
  • the image $P^{+}(Y_{n})$ of the positive-degree part of the reduced power subalgebra $P\subset A^{\sharp}$, and
  • all monomials $y\in Y_{n}$ such that $(\omega_{1}(y),\dots,\omega_{k}(y))<(n-1,\dots,n-1)$ in the lexicographic order.
He then uses an isomorphism $\psi:Y_{n}\to P_{n}$ sending $y_{i}$ to $x_{i}$ and intertwining the actions of $P$ and $A$ to identify $Y_{n}/G_{n}$ with the quotient $P_{n}/F_{n}$ from the classical hit problem in a suitable degree [3, Sec. 4].
On the $N_{n}$ side, Kameko defines a subspace $H_{n}\subset N_{n}$ spanned by $A^{\sharp}_{+}(N_{n})$ and those monomials $z$ for which
$(\omega_{0}(z),\omega_{1}(z),\dots,\omega_{k}(z))<(k,n-1,\dots,n-1)$
in the lexicographic order. He shows that the quotient $N_{n}/H_{n}$ decomposes as a direct sum over bigraded components $\Lambda^{a}_{n}\otimes(Y_{n}/G_{n})^{2b}$ with $a+2b=d$ and an explicit description of the contribution of $Q_{0}$ [3, Prop. 4.1].
We will only need the degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. In this case, Kameko constructs two sets of monomials $M_{0}$ and $M_{1}$ in $\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ and shows that their images in $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis. Moreover, he proves that
$M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\},$
and that the image of $Q_{0}$ in this summand is spanned, modulo $M_{0}$, by pairwise sums $\sigma_{1}(z_{k})+\sigma_{2}(z_{k})$; see [3, Props. 4.1, 5.1, 5.3].
We now turn to a more detailed analysis of this top summand.
3. A parity functional on Kameko’s top layer
In this section we work in the situation of the previous section with degree $d=k+2d_{1}$, $d_{1}=(n-1)(2^{k}-1)$, and monomial $z_{k}$ as in (2.2). We define a linear functional on the $\mathbb{F}_{2}$–span of the monomials in $M_{1}$ and show that its kernel exactly describes the intersection with the image of the motivic Steenrod algebra.
3.1. The $M_{1}$–summand and its parity functional
Fix integers $n\geq 2$ and $1\leq k<n$. Let $d_{1}=(n-1)(2^{k}-1)$ and $d=k+2d_{1}$. Following Kameko, we consider the summand
$\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}\subset(N_{n}/H_{n})_{d,*},$
and the sets $M_{0},M_{1}\subset\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ whose images form a basis of this summand. By [3, Eq. (4.3)], we have
(3.1) $M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\},$
and the images of $M_{0}\cup M_{1}$ in $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis.
Let $\pi:Y_{n}\to Y_{n}/G_{n}$ be the quotient map, and extend it to
$\pi:\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}\to\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Set
$V:=\Bigl\langle\pi\bigl(\sigma(z_{k})\bigr)\,\Big|\,\sigma\in\mathrm{Mono}(k)\Bigr\rangle\subset\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Since the $\pi(\sigma(z_{k}))$ are part of a basis, they are linearly independent and form a basis of $V$.
Definition 3.1 (Parity functional on the $M_{1}$–summand).
List $\mathrm{Mono}(k)$ as $\{\sigma_{1},\dots,\sigma_{N}\}$ and set $m_{i}=\sigma_{i}(z_{k})$, so $M_{1}=\{m_{1},\dots,m_{N}\}$ with $N=\binom{n}{k}$. Every element $v\in V$ can be written uniquely as
$v=\sum_{i=1}^{N}c_{i}\,\pi(m_{i}),\qquad c_{i}\in\mathbb{F}_{2}.$
We define
$\varepsilon:V\to\mathbb{F}_{2},\qquad\varepsilon(v)=\sum_{i=1}^{N}c_{i}\pmod{2}$.
Thus $\varepsilon(v)$ simply records whether the number of basis vectors $\pi(m_{i})$ that appear with coefficient $1$ in $v$ is even or odd.
3.2. A linear algebra lemma
We give a simple linear algebra fact that will be used repeatedly.
Lemma 3.2.
Let $W=\mathbb{F}_{2}^{N}$ with basis $e_{1},\dots,e_{N}$, and let $E\subset W$ be the subspace spanned by all pairwise sums $e_{i}+e_{j}$ with $i\neq j$. Then
$E=\{w=(w_{1},\dots,w_{N})\in W\mid w_{1}+\cdots+w_{N}=0\}.$
In other words, $E$ is exactly the hyperplane of vectors with even parity.
Proof.
For each $i\geq 2$ the vector $e_{1}+e_{i}$ lies in $E$. The N1N-1 vectors $e_{1}+e_{i}$ for $i=2,\dots,N$ are linearly independent and all have coordinate sum 0, so they span a subspace of dimension N1N-1 in the even-parity hyperplane. Thus $E$ has dimension at least N1N-1.
Conversely, let $w=(w_{1},\dots,w_{N})\in W$ satisfy $\sum_{i}w_{i}=0$. Then
$w=\sum_{i=1}^{N}w_{i}e_{i}=\sum_{i=2}^{N}w_{i}(e_{i}-e_{1})=\sum_{i=2}^{N}w_{i}(e_{1}+e_{i}),$
since over $\mathbb{F}_{2}$ we have $e_{1}=e_{1}$ and $w_{1}=\sum_{i=2}^{N}w_{i}$ by the parity assumption. Thus $w$ is a linear combination of the $e_{1}+e_{i}$, so $w\in E$.
We conclude that $E$ is an N1N-1–dimensional subspace of the even-parity hyperplane and coincides with it. ∎
3.3. The $M_{1}$–part of the $Q_{0}$–image
We now translate Kameko’s description of the image of $Q_{0}$ into a parity statement. Let $N_{n}=N_{n}(\mathbb{C})$ and $A^{\sharp}=A^{*,*}/(\tau)$ as before.
Proposition 3.3 (Kameko [3, Props. 4.1, 5.1, 5.3]).
Suppose $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ and $1\leq k<n$. Then:
  1. (1)
    The quotient $(N_{n}/H_{n})_{d,*}$ decomposes as
    $(N_{n}/H_{n})_{d,*}\cong\bigoplus_{a+2b=d}\Lambda^{a}_{n}\otimes(Y_{n}/G_{n})^{2b},$
    and the summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ has basis given by the images of the monomials in $M_{0}\cup M_{1}$.
  2. (2)
    For each monomial $z\in N_{n}$ of degree d1d-1 with $\omega_{0}(z)=k+1$, the image $\pi(Q_{0}(z))\in\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ is a linear combination of $\pi(z^{\prime})$ with $z^{\prime}\in M_{0}$ and elements of the form $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$ with $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$.
  3. (3)
    There exist monotone permutations $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ (explicitly, the ones with $\sigma_{1}(1)=1,\sigma_{1}(2)=3,\dots,\sigma_{1}(k)=k+1$ and $\sigma_{2}(1)=2,\dots,\sigma_{2}(k)=k+1$) and a monomial $z\in N_{n}^{d-1,*}$ with $(\omega_{0}(z),\omega_{1}(z))=(k+1,n-2)$ and $\alpha_{i}(z)=k$ for all $i$, such that the $M_{1}$–part of $\pi(Q_{0}(z))$ is precisely $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$.
The first two items are exactly Propositions 4.1 and 5.1 in [3]; the third is a reformulation of the second case in [3, Prop. 5.3] in our notation.
We now apply Lemma 3.2.
Proposition 3.4 (Parity of the $Q_{0}$–image in $V$).
With notation as above, identify $V$ with $W=\mathbb{F}_{2}^{N}$ by mapping the basis $\pi(\sigma_{i}(z_{k}))$ to $e_{i}$. Then the $M_{1}$–part of the image of
$Q_{0}:N_{n}^{d-1,*}\to N_{n}^{d,*}$
is exactly the even-parity hyperplane
$E=\{w\in W\mid w_{1}+\cdots+w_{N}=0\}\subset W.$
Equivalently,
$\varepsilon\bigl(\pi(Q_{0}(N_{n}^{d-1,*}))\cap V\bigr)=0.$
Proof.
By Proposition 3.3(2), the $M_{1}$–component of $\pi(Q_{0}(z))$ is always a linear combination of pairwise sums $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$, so the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ is contained in the subspace $E\subset W$ spanned by all $e_{i}+e_{j}$.
Conversely, by Proposition 3.3(3) there exist $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ and a monomial $z_{0}\in N_{n}^{d-1,*}$ such that the $M_{1}$–part of $\pi(Q_{0}(z_{0}))$ is precisely $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$. Let $i_{0},j_{0}$ be the indices corresponding to $\sigma_{1},\sigma_{2}$, so that this $M_{1}$–part is the vector $e_{i_{0}}+e_{j_{0}}$ in $W$.
Now let $S_{n}$ act on $N_{n}$ by permuting the variables $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$. This action preserves the subspaces $H_{n}$ and $G_{n}$, and it commutes with $Q_{0}$ and with the projection $\pi$, so for any permutation $\tau\in S_{n}$ we have
$\pi(Q_{0}(\tau\cdot z_{0}))=\tau\cdot\pi(Q_{0}(z_{0})).$
In particular, the $M_{1}$–part of $\pi(Q_{0}(\tau\cdot z_{0}))$ is $\pi(\tau\sigma_{1}(z_{k})+\tau\sigma_{2}(z_{k}))$, which corresponds to the vector $e_{i}+e_{j}$ where $i,j$ index the basis elements $\tau\sigma_{1}(z_{k})$ and $\tau\sigma_{2}(z_{k})$ of $V$.
As $\tau$ varies, the pairs of kk–subsets $\tau\sigma_{1}(\{1,\dots,k\})$ and $\tau\sigma_{2}(\{1,\dots,k\})$ run through all pairs of kk–subsets differing in exactly one element; that is, through all edges of the Johnson graph $J(n,k)$ on the vertex set $\mathrm{Mono}(k)$. It is well-known that the Johnson graph $J(n,k)$ is connected for $1\leq k\leq n-1,$ between any two basis elements $e_{i}$ and $e_{j}$ there is a path $i=v_{0},v_{1},\dots,v_{t}=j$ in $J(n,k)$, and in characteristic $2$ we have
$e_{i}+e_{j}=(e_{v_{0}}+e_{v_{1}})+(e_{v_{1}}+e_{v_{2}})+\cdots+(e_{v_{t-1}}+e_{v_{t}}).$
Each summand $e_{v_{\ell-1}}+e_{v_{\ell}}$ is realized as the $M_{1}$–part of some $\pi(Q_{0}(z))$ as above, so $e_{i}+e_{j}$ lies in the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$. Thus all vectors $e_{i}+e_{j}$ are contained in that image.
We conclude that the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ contains the subspace $E$ spanned by all $e_{i}+e_{j}$. Combined with the previous inclusion, this shows that it is exactly $E$.
By Lemma 3.2, $E$ is the even-parity hyperplane in $W$. In particular, $\varepsilon$ vanishes on $\pi(Q_{0}(N_{n}^{d-1,*}))\cap V$. ∎
3.4. The $A^{>0,*}$–image and the odd-parity gap
We next extend Proposition 3.4 from the image of $Q_{0}$ to the image of the full motivic Steenrod algebra in the summand $V$.
Lemma 3.5 (Reduced powers do not contribute to $M_{1}$).
In the summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ of $(N_{n}/H_{n})_{d,*}$, the $M_{1}$–component of the image of the reduced powers $\mathcal{P}^{a}$ ($a\geq 1$) is zero. Equivalently, modulo $H_{n}$ and modulo $\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$, the only contribution to $V$ from $A^{\sharp}_{+}(N_{n})$ comes from $Q_{0}$.
Proof.
By construction, $G_{n}\subset Y_{n}$ is spanned by $P^{+}(Y_{n})$ and monomials of smaller weight, where $P$ is the subalgebra of $A^{\sharp}$ generated by the reduced powers $\mathcal{P}^{a}$ [3, Sec. 4]. In particular, $P^{+}(Y_{n})\subset G_{n}$ by definition.
For $a\geq 1$, the operation $\mathcal{P}^{a}$ has bidegree $(2a,a)$ and acts trivially on $\Lambda_{n}$, so for any monomial $x\otimes y\in\Lambda^{k}_{n}\otimes Y_{n}^{2(d_{1}-a)}\subset N_{n}^{d-2a,*}$ we have
$\mathcal{P}^{a}(x\otimes y)=x\otimes\mathcal{P}^{a}(y)\in\Lambda^{k}_{n}\otimes\bigl(P^{+}(Y_{n})\cap Y_{n}^{2d_{1}}\bigr)\subset\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}.$
Here we used that $2(d_{1}-a)+2a=2d_{1}$, so the image lands in topological degree $d$ and polynomial degree $2d_{1}$.
The summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ of $(N_{n}/H_{n})_{d,*}$ is the image of $\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ under the quotient map $\pi:Y_{n}\to Y_{n}/G_{n}$. Since $\mathcal{P}^{a}(x\otimes y)\in\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$ for all $x\otimes y$, its image under $\pi$ is zero. Thus the $M_{1}$–component of the image of every $\mathcal{P}^{a}$ ($a\geq 1$) in this summand vanishes. Equivalently, modulo $H_{n}$ and $\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$, the only nontrivial image in $V$ of $A^{\sharp}_{+}(N_{n})$ comes from $Q_{0}$. ∎
Definition 3.6 (The $M_{1}$–projection to $V$).
Let $q:N_{n}\to N_{n}/H_{n}$ be the quotient map. In degree $d=k+2d_{1}$, Kameko’s decomposition [3, Prop. 4.1] gives a direct sum splitting of $(N_{n}/H_{n})_{d,*}$, hence a canonical projection
$\operatorname{pr}_{k,d_{1}}:(N_{n}/H_{n})_{d,*}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Let
$U_{0}:=\bigl\langle\pi(m)\mid m\in M_{0}\bigr\rangle\subset\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Since $\pi(M_{0})\cup\pi(M_{1})$ is a basis of $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$, we have a direct sum decomposition
$\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V,$
where $V=\langle\pi(M_{1})\rangle$ as in Section 3. Let
$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
be the projection with kernel $U_{0}$ (equivalently, $p_{M_{1}}$ kills $\pi(M_{0})$ and is the identity on $\pi(M_{1})$).
Finally, define the composite linear map
$\theta:N_{n}^{d,*}\xrightarrow{\ q\ }(N_{n}/H_{n})_{d,*}\xrightarrow{\ \operatorname{pr}_{k,d_{1}}\ }\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\xrightarrow{\ p_{M_{1}}\ }V.$
We call $\theta(u)$ the $M_{1}$–component (in $V$) of $u$.
Combining Proposition 3.4 and Lemma 3.5, we obtain the following.
Proposition 3.7 (Parity vanishes on the $A^{\sharp}_{+}$–image in the $M_{1}$–summand).
Let $u\in N_{n}^{d,*}$ be hit, i.e. $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$. Then the $M_{1}$–component $\theta(u)\in V$ has even parity:
$\varepsilon(\theta(u))=0.$
Equivalently,
$\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)\subseteq\ker(\varepsilon)\subset V.$
Proof.
By Lemma 3.5, in the summand $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ the reduced powers $\mathcal{P}^{a}$ ($a\geq 1$) contribute trivially to the $M_{1}$–part. Therefore, after applying the projection $\theta$, the only possible contribution from $A^{\sharp}_{+}(N_{n})$ comes from the Bockstein $Q_{0}$.
By Proposition 3.4, the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ is contained in the even-parity hyperplane $\ker(\varepsilon)\subset V$. Since $\theta$ is exactly the operation of passing to $(N_{n}/H_{n})_{d,*}$, projecting to the $(k,d_{1})$–summand, and then extracting the $M_{1}$–component, we obtain $\varepsilon(\theta(u))=0$ for every hit element $u$ of degree $d$. ∎
We can now state the main structural theorem of this section.
Theorem 3.8 (Theorem 1.4, General odd-parity non-hit theorem).
Let $n\geq 2$ and $1\leq k<n$, and let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. Let $M_{1}$ and $V$ be as in (3.1), and let $\varepsilon:V\to\mathbb{F}_{2}$ be the parity functional from Definition 3.1. Let $\theta:N_{n}^{d,*}\to V$ be the $M_{1}$–projection from Definition 3.6. Then:
  1. (1)
    One has an equality of subspaces of $V$:
    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
  2. (2)
    Consequently, $V/\theta(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*})$ is 11–dimensional over $\mathbb{F}_{2}$, generated by the class of any odd-parity element of $V$.
  3. (3)
    In particular, if $u\in\langle M_{1}\rangle\subset N_{n}^{d,*}$ has odd parity in its $M_{1}$–coordinates, then $u\notin A^{\sharp}_{+}(N_{n})$, so $u$ represents a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$.
Proof.
By Proposition 3.7, we have the inclusion
$\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)\subseteq\ker(\varepsilon).$
For the reverse inclusion, Proposition 3.4 shows that for any $i\neq j$ the vector $e_{i}+e_{j}$ (i.e. any pairwise sum $\pi(\sigma_{i}(z_{k})+\sigma_{j}(z_{k}))$) occurs as the $M_{1}$–part of $\pi(Q_{0}(z))$ for some $z\in N_{n}^{d-1,*}$, hence lies in $\theta(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*})$. These vectors span $\ker(\varepsilon)$ by Lemma 3.2. Therefore,
$\ker(\varepsilon)\subseteq\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr),$
which proves (1).
Since $\ker(\varepsilon)$ has codimension $1$ in $V$, statement (2) follows.
Finally, if $u\in\langle M_{1}\rangle$ has odd parity, then $\theta(u)=\pi(u)\in V$ also has odd parity (because $\theta$ restricts to the identity on $\pi(M_{1})$ by construction). If $u$ were hit, then $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$ would imply $\theta(u)\in\ker(\varepsilon)$ by (1), contradicting odd parity. Hence $u$ is not hit and defines a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$, and therefore in $QM_{n}^{d,*}(\mathbb{C})$. ∎
Corollary 3.9 (A large family of non-hit classes).
Let $S\subset\mathrm{Mono}(k)$ be a non-empty subset and set
$u_{S}:=\sum_{\sigma\in S}\sigma(z_{k})\in N_{n}^{d,*}.$
Let $\bar{u}_{S}$ be the image of $u_{S}$ in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$.
  1. (1)
    If $|S|$ is odd, then $\bar{u}_{S}$ is non-zero.
  2. (2)
    If $|S|$ is even, then the $M_{1}$–part of $\bar{u}_{S}$ lies in the image of $A^{\sharp}_{+}$, i.e. $\varepsilon(\bar{u}_{S})=0$.
In particular, each individual translate $\sigma(z_{k})$ represents a non-zero class, and any sum of an odd number of translates does as well.
Proof.
The $M_{1}$–component of $u_{S}$ is precisely $\sum_{\sigma\in S}\pi(\sigma(z_{k}))$, whose parity is $|S|$ modulo $2$. If $|S|$ is odd, this parity is $1$, so by Theorem 3.8 the class of $u_{S}$ is non-zero in the hit quotient. If $|S|$ is even, the parity is 0 and the $M_{1}$–component lies in $\ker(\varepsilon)$, which is the $M_{1}$–part of the $A^{\sharp}_{+}$–image. ∎
4. Arithmetic of the Kameko degrees
In this section we study the function $\beta(d)$ for the degrees $d=(n-1)(2^{k+1}-2)+k$ which arise in Kameko’s construction. The starting point is Proposition 2.1, which relates $\beta(d)>n$ to the inequality $\alpha(d+n)>n$.
4.1. Kameko’s original estimates
We recall Kameko’s computation from [3, Remark 2.2 and Proof of Prop. 1.5].
Proposition 4.1 (Kameko [3, Remark 2.2]).
Let $d=(n-1)(2^{k+1}-2)+k$.
  1. (1)
    If k=n1k=n-1, then $d+n=(n-1)\cdot 2^{n}+1$ and $\alpha(d+n)=\alpha(n-1)+1\leq n$.
  2. (2)
    If k=n2k=n-2, then $d+n=(n-1)\cdot 2^{n-1}$ and $\alpha(d+n)=\alpha(n-1)\leq n$.
Hence $\beta(d)\leq n$ for k=n1k=n-1 and k=n2k=n-2.
Proposition 4.2 (Kameko [3, Prop. 1.5]).
Let $d=(n-1)(2^{k+1}-2)+k$ with k=n3k=n-3. Then
$d+n=(n-1)\cdot 2^{n-2}-1=(n-2)\cdot 2^{n-2}+2^{n-3}+\cdots+2^{0}.$
If $\alpha(n-2)\geq 3$, then $\alpha(d+n)>n$, and hence $\beta(d)>n$.
Thus Kameko’s original family of motivic Peterson counterexamples arises from the case k=n3k=n-3. We now show that there are further families with k=n4k=n-4.
4.2. A new family with k=n4k=n-4
We continue to set $d=(n-1)(2^{k+1}-2)+k$ and $d_{1}=(n-1)(2^{k}-1)$, so that $d=k+2d_{1}$.
Lemma 4.3.
Let c=nkc=n-k. Then $1\leq c\leq n-1$ and
$d+n=(n-1)\cdot 2^{n-c+1}+(2-c).$
In particular, when $c=4$ (i.e. k=n4k=n-4) one has
$d+n=(n-1)\cdot 2^{n-3}-2.$
Proof.
By (2.3) and k=nck=n-c we have
$d$ $=(n-1)\bigl(2^{k+1}-2\bigr)+k$
$=(n-1)\bigl(2^{n-c+1}-2\bigr)+(n-c)$
$=(n-1)\cdot 2^{n-c+1}-2(n-1)+n-c.$
Adding $n$ gives
$d+n=(n-1)\cdot 2^{n-c+1}-2(n-1)+2n-c=(n-1)\cdot 2^{n-c+1}+(2-c),$
as claimed. For $c=4$ this simplifies to
$d+n=(n-1)\cdot 2^{n-3}-2.$
We now specialize to a particular infinite sequence of $n$.
Theorem 4.4 (Theorem 1.6, New numerical family with $\beta(d)>n$).
Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let $d=(n-1)(2^{k+1}-2)+k$ as above. Then $\beta(d)>n$.
Proof.
We use Proposition 2.1. Set c=nkc=n-k. Since k=n4k=n-4, we have $c=4$, so by Lemma 4.3 we obtain
$d+n=(n-1)\cdot 2^{n-3}-2.$
Now assume $n=2^{r}+1$ with $r\geq 1$, so $n-1=2^{r}$. Then
$d+n=2^{r}\cdot 2^{n-3}-2=2^{r+n-3}-2.$
Let L=r+n3L=r+n-3. The binary expansion of $2^{L}-1$ is a string of $L$ ones, so $2^{L}-2=(2^{L}-1)-1$ has binary expansion $11\cdots 10$ with L1L-1 ones. Thus
$\alpha(d+n)=\alpha(2^{L}-2)=L-1=(r+n-3)-1=r+n-4.$
Since $n=2^{r}+1$, we have
$\alpha(d+n)=r+2^{r}-3.$
We compare this with $n=2^{r}+1$:
$\alpha(d+n)>n\quad\Longleftrightarrow\quad r+2^{r}-3>2^{r}+1\quad\Longleftrightarrow\quad r-3>1\quad\Longleftrightarrow\quad r>4.$
By hypothesis $r\geq 5$, so indeed $\alpha(d+n)>n$. By Proposition 2.1, this is equivalent to $\beta(d)>n$. ∎
Remark 4.5.
For small values of $r$ one can verify the above calculation by direct computation. For example, when $r=5$ we have $n=33$, $k=29$ and $d+n=2^{35}-2$, whose binary expansion has $34$ ones, so $\alpha(d+n)=34>33=n$. For $r=6$ we obtain $n=65$, $k=61$ and $d+n=2^{68}-2$ with $\alpha(d+n)=67>65$, and so on.
Theorem 4.4 gives an infinite family of pairs $(n,k)$ with k=n4k=n-4 for which $\beta(d)>n$. In the next section we combine this with the odd-parity non-hit theorem to produce new motivic Peterson counterexamples.
5. Base change and motivic counterexamples over general fields
So far we have worked with the base field $\mathbb{C}$, as in Kameko’s original paper. We now explain why the constructions and counterexamples extend to any algebraically closed field $K$ of characteristic 0.
5.1. Motivic cohomology and the Steenrod algebra over general base fields
Let $K$ be an algebraically closed field of characteristic 0. Voevodsky constructed the mod $2$ motivic Steenrod algebra $A_{K}^{*,*}$ over any field of characteristic different from $2$ and described it in terms of generators and relations that depend only on the characteristic [8]. In particular, if $K$ and $\mathbb{C}$ are algebraically closed of characteristic 0, then $A_{K}^{*,*}\cong A_{\mathbb{C}}^{*,*}$ as bigraded Hopf algebras.
Similarly, for an elementary abelian 22–group $V_{n}=(\mathbb{Z}/2)^{n}$, the motivic cohomology ring $H^{*,*}(BV_{n}/K;\mathbb{F}_{2})$ is computed in terms of the cohomology of $B\mu_{2}$ and is independent of the choice of algebraically closed field of characteristic 0. Concretely, one has an isomorphism
(5.1) $M_{n}(K):=H^{*,*}(BV_{n}/K;\mathbb{F}_{2})\cong\mathbb{F}_{2}[\tau,x_{1},\dots,x_{n},y_{1},\dots,y_{n}]/(x_{1}^{2}+\tau y_{1},\dots,x_{n}^{2}+\tau y_{n}),$
with $|\tau|=(0,1)$, $|x_{i}|=(1,1)$ and $|y_{i}|=(2,1)$, and the action of $A_{K}^{*,*}$ is given by the same formulas as over $\mathbb{C}$; see, for example, [8, Sec. 9] together with [2]. In particular, there is a natural isomorphism of bigraded $A^{*,*}$–modules between $M_{n}(K)$ and $M_{n}(\mathbb{C})$, and likewise for their quotients $N_{n}(K)=M_{n}(K)/(\tau)$ and $N_{n}(\mathbb{C})$.
Remark 5.1.
No field embedding $K\hookrightarrow\mathbb{C}$ is needed here. For algebraically closed fields of characteristic 0, one has $\rho=[-1]=0\in H^{1,1}(\operatorname{Spec}K;\mathbb{F}_{2})$, just as for $K=\mathbb{C}$. In particular, the presentations (2.1) and (5.1) have the same defining relations, and the motivic Steenrod operations $Q_{0},\mathcal{P}^{a}$ act by the same generator formulas (together with Cartan) on the generators $\tau,x_{i},y_{i}$. Hence, using these explicit presentations, the assignment
$\tau\mapsto\tau,\qquad x_{i}\mapsto x_{i},\qquad y_{i}\mapsto y_{i}$
defines an isomorphism of bigraded rings $\Phi_{K}:M_{n}(K)\xrightarrow{\cong}M_{n}(\mathbb{C})$, and because the Steenrod action formulas on generators coincide, $\Phi_{K}$ is in fact an isomorphism of bigraded $A^{*,*}$–modules.
Using (5.1), we define the motivic hit quotients
$QM_{n}^{d,*}(K):=M_{n}(K)_{d,*}/(A^{>0,*}\cdot M_{n}(K)\cap M_{n}(K)_{d,*})$
exactly as in the $\mathbb{C}$–case, and similarly for the quotients $QN_{n}^{d,*}(K)$ of $N_{n}(K)$. The formulas and decompositions from Sections 2 and 3 make sense over $K$ and are preserved under the isomorphism $M_{n}(K)\cong M_{n}(\mathbb{C})$.
5.2. Extension of Kameko’s non-hit classes
We now show that Kameko’s non-hit monomials and our parity refinements extend to arbitrary algebraically closed fields of characteristic 0.
Proposition 5.2.
Let $K$ be an algebraically closed field of characteristic 0. For any integers $n\geq 1$ and $1\leq k<n$, the monomial $z_{k}\in M_{n}(K)$ defined by (2.2) is not in $A^{>0,*}\cdot M_{n}(K)$.
Proof.
By Remark 5.1, there is an isomorphism of bigraded $A^{*,*}$–modules
$\Phi_{K}:M_{n}(K)\xrightarrow{\ \cong\ }M_{n}(\mathbb{C})$
sending $\tau,x_{i},y_{i}$ to the corresponding generators. In particular, $\Phi_{K}(z_{k})$ is the monomial $z_{k}$ in $M_{n}(\mathbb{C})$ considered by Kameko.
Suppose $z_{k}\in A^{>0,*}\cdot M_{n}(K)$. Applying $\Phi_{K}$, we obtain $\Phi_{K}(z_{k})\in A^{>0,*}\cdot M_{n}(\mathbb{C})$. But by Proposition 2.3, $\Phi(z_{k})$ is not in this image. This is a contradiction. Hence $z_{k}$ is not hit in $M_{n}(K)$. ∎
The same argument applies to any linear combination of the translates $\sigma(z_{k})$.
Proposition 5.3.
Let $K$ be an algebraically closed field of characteristic 0, and let $n,k,d,M_{1},V,\varepsilon$ be as in Theorem 3.8. Then the conclusions of Theorem 3.8 and Corollary 3.9 hold over $K$: in particular, any odd-parity linear combination of the $\sigma(z_{k})$ represents a non-zero class in $QM_{n}^{d,*}(K)$.
Proof.
By Remark 5.1 there is an isomorphism of bigraded $A^{*,*}$–modules
$\Phi_{K}:M_{n}(K)\xrightarrow{\ \cong\ }M_{n}(\mathbb{C})$
sending $\tau,x_{i},y_{i}$ to the corresponding generators. Since $\Phi_{K}(\tau)=\tau$, it induces an isomorphism of $A^{\sharp}$–modules
$\Phi_{K}^{\sharp}:N_{n}(K)=M_{n}(K)/(\tau)\xrightarrow{\ \cong\ }M_{n}(\mathbb{C})/(\tau)=N_{n}(\mathbb{C}).$
Passing to hit quotients, we obtain an induced isomorphism
$\overline{\Phi}_{K}:QN_{n}^{d,*}(K)\xrightarrow{\ \cong\ }QN_{n}^{d,*}(\mathbb{C}).$
Under these identifications, the definitions of the subspaces $H_{n}$, $G_{n}$ and the sets $M_{0},M_{1}$ (hence also the subspace $V$ and the parity functional $\varepsilon$) are preserved. Moreover, the $M_{1}$–projection $\theta$ from Definition 3.6 is functorial with respect to $\Phi_{K}^{\sharp}$ (it is defined using the same direct-sum splitting and the same basis $\pi(M_{0})\cup\pi(M_{1})$), hence $\overline{\Phi}_{K}$ carries odd-parity classes to odd-parity classes.
Therefore, if $u=\sum_{\sigma\in S}\sigma(z_{k})$ has odd parity, then its image $\overline{\Phi}_{K}(\bar{u})\in QN_{n}^{d,*}(\mathbb{C})$ is an odd-parity class, which is non-zero by Theorem 3.8 (over $\mathbb{C}$). Since $\overline{\Phi}_{K}$ is an isomorphism, $\bar{u}\neq 0$ in $QN_{n}^{d,*}(K)$. Finally, using $QN_{n}^{d,*}(K)\cong QM_{n}^{d,*}(K)/\tau\,QM_{n}^{d,*}(K)$, we conclude that $u$ represents a non-zero class in $QM_{n}^{d,*}(K)$ as well. ∎
5.3. Motivic Peterson counterexamples over $K$
We now combine the arithmetic from Section 4 with the parity statement over general base fields.
Theorem 5.4 (Motivic Peterson counterexamples for k=n3k=n-3 over $K$).
Let $K$ be an algebraically closed field of characteristic 0. Let $n\geq 9$ be an integer with $\alpha(n-2)\geq 3$, set k=n3k=n-3, and let $d=(n-1)(2^{k+1}-2)+k$. Then
  1. (1)
    $\beta(d)>n$.
  2. (2)
    The motivic hit quotient $QM_{n}^{d,*}(K)$ is non-zero.
In particular, Kameko’s motivic Peterson-type conjecture fails in these degrees over every such $K$.
Proof.
The inequality $\beta(d)>n$ is purely numerical and independent of the base field; it is proved in Proposition 4.2.
For the non-vanishing of $QM_{n}^{d,*}(K)$, we may argue as follows. Over $\mathbb{C}$, Theorem 1.3 shows that $z_{k}$ is non-hit and has degree $d$. Hence $QM_{n}^{d,*}(\mathbb{C})\neq 0$. The isomorphism $M_{n}(K)\cong M_{n}(\mathbb{C})$ of $A^{*,*}$–modules induces an isomorphism
$QM_{n}^{d,*}(K)\cong QM_{n}^{d,*}(\mathbb{C}),$
so $QM_{n}^{d,*}(K)\neq 0$ as well. Alternatively, one can invoke Proposition 5.2, which implies that $z_{k}$ represents a non-zero class in $QM_{n}^{d,*}(K)$. ∎
We now state and prove the analogue for the new family of degrees from Theorem 4.4.
Theorem 5.5 (New motivic Peterson counterexamples at distance $4$).
Let $K$ be an algebraically closed field of characteristic 0. Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let $d=(n-1)(2^{k+1}-2)+k$ be the degree of $z_{k}$. Then:
  1. (1)
    $\beta(d)>n$.
  2. (2)
    The motivic hit quotient $QM_{n}^{d,*}(K)$ is non-zero.
  3. (3)
    More precisely, for any non-empty subset $S\subset\mathrm{Mono}(k)$ with $|S|$ odd, the element
    $u_{S}=\sum_{\sigma\in S}\sigma(z_{k})\in M_{n}(K)_{d,*}$
    represents a non-zero class in $QM_{n}^{d,*}(K)$.
Thus these degrees $(n,d)$ form an infinite family of counterexamples to the motivic Peterson-type conjecture over $K$, distinct from the family with k=n3k=n-3.
Proof.
The inequality $\beta(d)>n$ follows from Theorem 4.4 and Proposition 2.1; this is purely number-theoretic and independent of $K$.
For the non-vanishing of $QM_{n}^{d,*}(K)$ and the explicit non-hit elements, we use Proposition 5.3. In the present situation, the degree $d$ is again of the form $k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, so the odd-parity non-hit theorem applies. In particular, any $u_{S}$ with $|S|$ odd has odd parity in the $M_{1}$–summand and hence defines a non-zero class in $QN_{n}^{d,*}(K)$ and therefore in $QM_{n}^{d,*}(K)$.
In particular, taking $S=\{\mathrm{id}\}$ shows that $z_{k}$ itself is non-hit and yields $QM_{n}^{d,*}(K)\neq 0$. This completes the proof. ∎
Combining Theorems 5.4 and 5.5 proves Theorem 1.7 from the introduction.
Remark 5.6.
The family in Theorem 5.5 is genuinely distinct from Kameko’s original family [3]. In Kameko’s family one has k=n3k=n-3, while here k=n4k=n-4; the corresponding degrees $d$ differ, and the two families occupy different regions of the (n,d)(n,d)–plane.

### v3

@@
 Abstract.
-The motivic hit problem seeks a minimal generating set for the motivic cohomology of classifying spaces as a module over the motivic Steenrod algebra.
-In this paper, we provide a complete structural classification of the hit image within the top layer of Kameko’s decomposition, revealing a sharp dichotomy governed by parity.
-Specifically, for degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, let $V$ denote the span of the monotone translates of Kameko’s monomial $z_{k}$.
-We prove that the intersection of $V$ with the hit subspace is exactly the even-parity hyperplane.
-Consequently, the quotient $V/(\text{hits})$ is one-dimensional and is generated by any odd-parity linear combination of these translates.
-This establishes the ”odd-parity gap” as the precise obstruction to the hit property in this layer.
-As an arithmetic consequence of this structural result, we identify a new infinite numerical family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$) where $\beta(d)>n$.
-This yields a new class of counterexamples to the motivic Peterson-type conjecture, distinct from Kameko’s original family.
-Furthermore, we show that these structural results and counterexamples persist over any algebraically closed field of characteristic 0.
+The motivic hit problem asks for a minimal set of generators of $H^{*,*}(BV_{n};\mathbb{F}_{2})$ as a module over the motivic Steenrod algebra.
+Masaki Kameko [3] constructed, in a distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a decomposition in which the top layer is spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates.
+While Kameko obtained a restriction on the possible $Q_{0}$–image in the top layer, we determine the corresponding hit image in $V$ exactly.
+We make three main contributions as follows:
+First, we give an exact structural description of the hit subspace on the top-layer summand $V=\langle\sigma(z_{k})\rangle$ by constructing a parity functional $\varepsilon:V\to\mathbb{F}_{2}$ and proving that $V\cap(\text{hit subspace})=\ker(\varepsilon)$; equivalently, parity yields a complete criterion for hitness in the $M_{1}$–summand and $V/(\text{hits})\cong\mathbb{F}_{2}$.
+In particular, this computes the quotient $V/(\text{hits})$.
+We next explain how this parity classification interacts with the numerical condition $\beta(d)>n$.
+In particular, the counterexamples arise as consequences of the structural theorem.
+Second, we combine this classification with the numerical condition $\beta(d)>n$ to obtain a new infinite counterexample family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$), distinct from Kameko’s k=n3k=n-3 family.
+Third, we prove a base-change invariance statement: the parity classification on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
 Key words and phrases:
@@
 Moreover, the condition is also necessary in the sense that the monomial $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ never lies in $A^{+}(P_{n})$.
-The interplay between the hit problem and the cohomology of $A$ via Singer’s transfer [7] connects this story to the classical Adams spectral sequence and the stable homotopy groups of spheres; see, for example, Ault [1], Pengelley–Williams [4], the present author [5, 6].
+The interplay between the hit problem and the cohomology of $A$ via Singer’s transfer [5] connects this story to the classical Adams spectral sequence and the stable homotopy groups of spheres; see, e.g.
+[1], [4], [7, 8].
 In motivic homotopy theory, Voevodsky introduced a motivic Steenrod algebra $A^{*,*}$ and reduced power operations in motivic cohomology over a base field of characteristic different from $2$ [6].
@@
 He then shows that only certain linear combinations can appear in the image of the Bockstein $Q_{0}$, and that $z_{k}$ itself is not among them.
-The main objective of this paper is to uncover the sharp structural dichotomy governing the top layer of the motivic hit problem.
-While previous work focused on the existence of counterexamples, we provide a complete linear-algebraic characterization of the hit image in terms of parity.
+The present paper strengthens Kameko’s analysis of the top layer from an inclusion statement to an exact structural description. In Kameko’s decomposition, the summand
+$V=\big\langle\sigma(z_{k})\ \big|\ \sigma\in\mathrm{Mono}(k)\big\rangle$
+captures the entire “top layer” contribution in degree $d=k+2d_{1}$.
+Kameko showed that only certain linear combinations can arise from the Bockstein $Q_{0}$ in this layer, namely combinations generated by pairwise sums $\sigma_{1}(z_{k})+\sigma_{2}(z_{k})$, yielding a containment for the possible $Q_{0}$–image in $V$.
+What remained open is whether this containment is sharp, i.e.
+whether it actually exhausts the hit image inside $V$, and hence what the precise obstruction to being hit is.
+We answer this question by giving a complete parity-based description of the hit image in $V$.
+Accordingly, our main contribution in this work is structural: we upgrade Kameko’s top-layer restriction from a containment to an equality by giving an exact parity-based description of the hit image inside $V$.
+In particular, the counterexample families arise as consequences of this classification, once combined with arithmetic input.
   • (1)
-    Structural Classification: We isolate a parity functional on the top layer of Kameko’s decomposition and prove that the hit image is exactly the even-parity hyperplane.
-This implies that the obstruction to the hit problem in this layer is one-dimensional and is completely determined by parity.
+    Parity classification (exact hit image).
+We construct an explicit linear functional $\varepsilon:V\to\mathbb{F}_{2}$ (the parity) and prove an if-and-only-if criterion: a class in $V$ is hit precisely when it has even parity.
+Equivalently, the hit image in $V$ is the codimension-one hyperplane $\ker(\varepsilon)$, so that $V/(\text{hits})\cong\mathbb{F}_{2}$.
   • (2)
-    Arithmetic Consequences: We apply this structural insight to the arithmetic of degrees $d=(n-1)(2^{k+1}-2)+k$.
+    Arithmetic application (new counterexample family).
+As a consequence, once $\beta(d)>n$ holds in these degrees, any odd-parity class in $V$ yields a counterexample to the motivic Peterson-type conjecture.
 Using the equivalence between $\beta(d)>n$ and $\alpha(d+n)>n$ [3, Prop.
-2.1], we identify a new infinite family of pairs $(n,k)$ with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) where the hit quotient is non-zero despite $\beta(d)>n$.
+2.1], we establish a new infinite family with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) for which $\beta(d)>n$.
   • (3)
-    Generalization: We demonstrate that these results are not specific to $\mathbb{C}$ but persist over any algebraically closed field $K$ of characteristic 0, relying on the stability of the motivic Steenrod algebra under base change.
+    Base change.
+We show that the parity classification on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
 Our main structural result is the following, which establishes that parity is the precise obstruction to being hit in the summand $V$.
-Theorem 1.4 (Structure of the Top Hit Layer).
-For any integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ be the degree of Kameko’s monomial $z_{k}$.
-Let $M_{1}$ be the set of all translates $\sigma(z_{k})$ under monotone permutations $\sigma\in\mathrm{Mono}(k)$, and let $V$ be their span in the motivic hit quotient $QN_{n}^{d,*}$.
-Let $\theta:N_{n}^{d,*}\to V$ be the map extracting the $M_{1}$–component (defined precisely in Section 3). Then there is a linear functional
-$\varepsilon:V\to\mathbb{F}_{2}$
-such that:
-  1. (1)
-    Exact Image Characterization: One has
-    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon)\subset V.$
-  2. (2)
-    Non-hit Criterion: Consequently, any class whose $M_{1}$–component has odd parity is non-hit. The quotient space V/(hit image)V/($\text${hit image}) is exactly one-dimensional.
-Remark 1.5 (Minimality of the Obstruction).
-Theorem 1.4 shows that the description of the hit image as the even-parity hyperplane is sharp.
-The “odd-parity gap” is the smallest possible non-trivial quotient, implying that no further reduction by the Steenrod algebra is possible in this layer.
+Theorem 1.4 (Parity classification in the top layer).
+For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span in $QN_{n}^{d,*}$ of the monotone translates $\sigma(z_{k})$ with $\sigma\in\mathrm{Mono}(k)$.
+Let $\theta:N_{n}^{d,*}\to V$ be the projection onto the $M_{1}$–component (defined in Section 3).
+There exists a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
+  1. (i)
+    Exact hit image in $V$:
+    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
+  2. (ii)
+    If and only if criterion: For $v\in V$, one has $v$ is hit in $QN_{n}^{d,*}$ if and only if $\varepsilon(v)=0$.
+In particular $V/(\text{hits})\cong\mathbb{F}_{2}$, generated by any $v\in V$ with $\varepsilon(v)=1$.
+Remark 1.5 (Sharpness).
+Theorem 1.4 upgrades Kameko’s restriction [3] on the possible $Q_{0}$–image in the top layer to an equality statement: the hit image inside $V$ is a single explicit codimension-one hyperplane.
+Thus the obstruction in this summand is minimal (one-dimensional) and is completely detected by parity.
+Equivalently, the theorem computes the one-dimensional quotient $V/(\text{hits})$, and provides an effective parity criterion: a class in $V$ is hit if and only if it has even parity.
 On the arithmetic side, we prove:
@@
 (5.1) $M_{n}(K):=H^{*,*}(BV_{n}/K;\mathbb{F}_{2})\cong\mathbb{F}_{2}[\tau,x_{1},\dots,x_{n},y_{1},\dots,y_{n}]/(x_{1}^{2}+\tau y_{1},\dots,x_{n}^{2}+\tau y_{n}),$
-with $|\tau|=(0,1)$, $|x_{i}|=(1,1)$ and $|y_{i}|=(2,1)$, and the action of $A_{K}^{*,*}$ is given by the same formulas as over $\mathbb{C}$; see, for example, [8, Sec.
+with $|\tau|=(0,1)$, $|x_{i}|=(1,1)$ and $|y_{i}|=(2,1)$, and the action of $A_{K}^{*,*}$ is given by the same formulas as over $\mathbb{C}$; see, for example, [6, Sec.
 9] together with [2].

### v4

@@
 The motivic hit problem asks for a minimal set of generators of $H^{*,*}(BV_{n};\mathbb{F}_{2})$ as a module over the motivic Steenrod algebra.
-Masaki Kameko [3] constructed, in a distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a decomposition in which the top layer is spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates.
-While Kameko obtained a restriction on the possible $Q_{0}$–image in the top layer, we determine the corresponding hit image in $V$ exactly.
-We make three main contributions as follows:
-First, we give an exact structural description of the hit subspace on the top-layer summand $V=\langle\sigma(z_{k})\rangle$ by constructing a parity functional $\varepsilon:V\to\mathbb{F}_{2}$ and proving that $V\cap(\text{hit subspace})=\ker(\varepsilon)$; equivalently, parity yields a complete criterion for hitness in the $M_{1}$–summand and $V/(\text{hits})\cong\mathbb{F}_{2}$.
-In particular, this computes the quotient $V/(\text{hits})$.
-We next explain how this parity classification interacts with the numerical condition $\beta(d)>n$.
-In particular, the counterexamples arise as consequences of the structural theorem.
-Second, we combine this classification with the numerical condition $\beta(d)>n$ to obtain a new infinite counterexample family with $n=2^{r}+1$ and k=n4k=n-4 ($r\geq 5$), distinct from Kameko’s k=n3k=n-3 family.
-Third, we prove a base-change invariance statement: the parity classification on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
+Masaki Kameko [3] constructed, in the distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a top layer spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates.
+In this paper, we work on the raw degree–$d$ component $N_{n}^{d,*}$, before passing to the quotient by the hit subspace.
+We therefore construct a local projection
+$\vartheta:N_{n}^{d,*}\longrightarrow V,$
+from the raw degree–$d$ component to Kameko’s $M_{1}$–summand $V=\langle\{\sigma(z_{k})|\ \sigma\in\mathrm{Mono}(k)\}\rangle$, together with a parity functional $\varepsilon:V\to\mathbb{F}_{2}$, and prove that
+$\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
+Thus parity gives an exact description of the local top-layer image of hit elements.
+As a consequence, any element whose local top-layer component has odd parity is non-hit.
+In particular, every odd-parity linear combination of the translates $\sigma(z_{k})$ represents a non-zero class in the motivic hit quotient.
+Subsequently, a direct binary calculation shows that for
+$n=2^{r}+1,\qquad k=n-4,\qquad r\geq 5,$
+one has $\beta(d)>n$ for
+$d=(n-1)(2^{k+1}-2)+k.$
+Combined with Kameko’s non-hit theorem, this yields a new infinite family of counterexamples to the motivic Peterson-type conjecture, distinct from Kameko’s k=n3k=n-3 family.
+Our local parity criterion sharpens this conclusion by showing that every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit in those degrees.
+Finally, we prove a base-change invariance statement: the local parity criterion on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
 Key words and phrases:
@@
 Kameko’s proof is quite delicate.
-He passes to the quotient $N_{n}(\mathbb{C})=M_{n}(\mathbb{C})/(\tau)$ and constructs a family of monomials in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ that give a basis for the top layer of the motivic hit quotient.
+He passes to the quotient $N_{n}(\mathbb{C})=M_{n}(\mathbb{C})/(\tau)$ and constructs a family of monomials in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ that give a basis for the top layer in the relevant filtration.
 He then shows that only certain linear combinations can appear in the image of the Bockstein $Q_{0}$, and that $z_{k}$ itself is not among them.
-The present paper strengthens Kameko’s analysis of the top layer from an inclusion statement to an exact structural description. In Kameko’s decomposition, the summand
-$V=\big\langle\sigma(z_{k})\ \big|\ \sigma\in\mathrm{Mono}(k)\big\rangle$
-captures the entire “top layer” contribution in degree $d=k+2d_{1}$.
-Kameko showed that only certain linear combinations can arise from the Bockstein $Q_{0}$ in this layer, namely combinations generated by pairwise sums $\sigma_{1}(z_{k})+\sigma_{2}(z_{k})$, yielding a containment for the possible $Q_{0}$–image in $V$.
-What remained open is whether this containment is sharp, i.e.
-whether it actually exhausts the hit image inside $V$, and hence what the precise obstruction to being hit is.
-We answer this question by giving a complete parity-based description of the hit image in $V$.
-Accordingly, our main contribution in this work is structural: we upgrade Kameko’s top-layer restriction from a containment to an equality by giving an exact parity-based description of the hit image inside $V$.
-In particular, the counterexample families arise as consequences of this classification, once combined with arithmetic input.
+The present paper strengthens Kameko’s top-layer analysis by working on the raw degree–$d$ component $N_{n}^{d,*}$, before passing to the quotient by the full hit subspace.
+The key point is that, in degree $d=k+2d_{1}$, the part detected by Kameko’s set
+$V=\bigl\langle\pi(\sigma(z_{k}))\ \big|\ \sigma\in\mathrm{Mono}(k)\bigr\rangle\subset\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
+should be viewed as a local top layer, rather than as a subspace of the final quotient by hits. Here and throughout, $\mathrm{Mono}(k)$ denotes the set of monotone injections
+$\sigma:\{1,\dots,k\}\hookrightarrow\{1,\dots,n\}.$
+Equivalently, $\mathrm{Mono}(k)$ may be identified with the set of kk–subsets of $\{1,\dots,n\}$ via $\sigma\mapsto\mathrm{Im}(\sigma)$.
+Once this distinction is made, Kameko’s containment [3] for the $Q_{0}$–image becomes an exact parity statement for the local image of hit elements.
+Our main contribution is therefore structural: we construct a local projection
+$\vartheta:N_{n}^{d,*}\longrightarrow V$
+and an explicit parity functional $\varepsilon:V\to\mathbb{F}_{2}$ such that
+$\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
+This gives a complete description of which vectors occur as the $M_{1}$–part of hit elements, and it yields an effective odd-parity obstruction.
   • (1)
-    Parity classification (exact hit image).
-We construct an explicit linear functional $\varepsilon:V\to\mathbb{F}_{2}$ (the parity) and prove an if-and-only-if criterion: a class in $V$ is hit precisely when it has even parity.
-Equivalently, the hit image in $V$ is the codimension-one hyperplane $\ker(\varepsilon)$, so that $V/(\text{hits})\cong\mathbb{F}_{2}$.
+    Local parity criterion.
+We prove that, in degree $d=k+2d_{1}$, the local top-layer image of hit elements is exactly the codimension-one hyperplane $\ker(\varepsilon)\subset V$.
+Equivalently, an element of $N_{n}^{d,*}$ whose local top-layer component has odd parity cannot be hit.
+In particular, every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit.
   • (2)
-    Arithmetic application (new counterexample family).
-As a consequence, once $\beta(d)>n$ holds in these degrees, any odd-parity class in $V$ yields a counterexample to the motivic Peterson-type conjecture.
-Using the equivalence between $\beta(d)>n$ and $\alpha(d+n)>n$ [3, Prop.
-2.1], we establish a new infinite family with k=n4k=n-4 (distinct from Kameko’s k=n3k=n-3 case) for which $\beta(d)>n$.
+    Arithmetic consequence and refinement. A direct binary calculation shows that if $n=2^{r}+1$ and k=n4k=n-4 with $r\geq 5$, then $\beta(d)>n$ for
+    $d=(n-1)(2^{k+1}-2)+k.$
+    Combined with Kameko’s non-hit theorem [3], this gives a new infinite family of motivic Peterson counterexamples.
+The local parity criterion refines this by proving that every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit in those degrees.
   • (3)
     Base change.
-We show that the parity classification on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
-Our main structural result is the following, which establishes that parity is the precise obstruction to being hit in the summand $V$.
-Theorem 1.4 (Parity classification in the top layer).
-For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span in $QN_{n}^{d,*}$ of the monotone translates $\sigma(z_{k})$ with $\sigma\in\mathrm{Mono}(k)$.
-Let $\theta:N_{n}^{d,*}\to V$ be the projection onto the $M_{1}$–component (defined in Section 3).
-There exists a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
+We show that the local parity criterion on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
+Our main structural result is the following.
+Theorem 1.4 (Local parity criterion in the top layer).
+For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span of the monotone translates $\sigma(z_{k})$ in Kameko’s $M_{1}$–summand.
+There exist a linear map
+$\vartheta:N_{n}^{d,*}\longrightarrow V$
+and a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
   1. (i)
-    Exact hit image in $V$:
-    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
+    Exact local top-layer hit image:
+    $\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
   2. (ii)
-    If and only if criterion: For $v\in V$, one has $v$ is hit in $QN_{n}^{d,*}$ if and only if $\varepsilon(v)=0$.
-In particular $V/(\text{hits})\cong\mathbb{F}_{2}$, generated by any $v\in V$ with $\varepsilon(v)=1$.
-Remark 1.5 (Sharpness).
-Theorem 1.4 upgrades Kameko’s restriction [3] on the possible $Q_{0}$–image in the top layer to an equality statement: the hit image inside $V$ is a single explicit codimension-one hyperplane.
-Thus the obstruction in this summand is minimal (one-dimensional) and is completely detected by parity.
-Equivalently, the theorem computes the one-dimensional quotient $V/(\text{hits})$, and provides an effective parity criterion: a class in $V$ is hit if and only if it has even parity.
-On the arithmetic side, we prove:
-Theorem 1.6 (New numerical family with $\beta(d)>n$).
-Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let
-$d=(n-1)\bigl(2^{k+1}-2\bigr)+k$
-be the degree of $z_{k}$. Then $\beta(d)>n$.
-Combining Theorems 1.3, 1.4 and 1.6 with the base-change discussion in Section 5, we obtain the following.
+    Odd-parity obstruction: if $u\in N_{n}^{d,*}$ satisfies $\varepsilon(\vartheta(u))=1$, then $u\notin A_{+}^{\sharp}(N_{n})$.
+  3. (iii)
+    In particular, any odd-parity linear combination of the monomials $\sigma(z_{k})$ is non-hit and therefore represents a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$ and in $QM_{n}^{d,*}(\mathbb{C})$.
+Remark 1.5 (Why the local formulation matters).
+The map $\vartheta$ is defined on the raw degree–$d$ piece $N_{n}^{d,*}$, before quotienting by $A_{+}^{\sharp}(N_{n})$.
+This is essential: if one first passes to a quotient that already modds out hit elements, then every hit class becomes zero and the top-layer information is lost.
+The theorem above isolates the correct statement: parity detects exactly the local top-layer parts of hit elements.
+On the arithmetic side, we show by a direct binary calculation that for $r\geq 5$, the choice $n=2^{r}+1$ and k=n4k=n-4 satisfies $\beta(d)>n$ for
+$d=(n-1)(2^{k+1}-2)+k.$
+Together with Kameko’s non-hit theorem, this yields a new infinite family of motivic Peterson counterexamples, distinct from Kameko’s original case k=n3k=n-3.
+The local parity criterion strengthens this conclusion by showing that every odd-parity linear combination of the translates $\sigma(z_{k})$ represents a non-zero class in the corresponding motivic hit quotient.
+Combining Theorem 1.3, the local parity criterion of Theorem 1.4, the direct binary calculation for the case $n=2^{r}+1$, k=n4k=n-4, and the base-change discussion in Section 4, we obtain the following.
 Theorem 1.6.
@@
   3. (3)
-    More precisely, any odd-parity linear combination of the monomials $\sigma(z_{k})\in M_{n}^{d,*}(K)$ with $\sigma\in\mathrm{Mono}(k)$ represents a non-zero class in $QM_{n}^{d,*}(K)$.
-Thus, combining the parity description of the top-layer hit image with the numerical condition $\beta(d)>n$, we obtain an infinite family of motivic Peterson-type counterexamples over any algebraically closed field of characteristic 0, distinct from Kameko’s k=n3k=n-3 family [3].
+    More precisely, any odd-parity linear combination of the monomials $\sigma(z_{k})\in M_{n}(K)_{d,*}$, with $\sigma\in\mathrm{Mono}(k)$, represents a non-zero class in $QM_{n}^{d,*}(K)$.
+Thus, combining the local parity description of the top layer with the direct verification of the inequality $\beta(d)>n$ in the case $n=2^{r}+1$, k=n4k=n-4, we obtain an infinite family of motivic Peterson-type counterexamples over any algebraically closed field of characteristic 0, distinct from Kameko’s k=n3k=n-3 family [3].
 The paper is organized as follows.
 In Section 2 we review the classical hit problem, the motivic Steenrod algebra, and Kameko’s description of the motivic cohomology of $BV_{n}$ and the motivic hit problem.
-Section 3 introduces the parity functional on the top layer of the motivic hit quotient and proves Theorem 1.4.
-Section 4 is devoted to the arithmetic of the degrees $d=(n-1)(2^{k+1}-2)+k$, including Kameko’s original estimates and the new family in Theorem 1.6.
-Finally, Section 5 discusses the extension to arbitrary algebraically closed base fields of characteristic 0 and proves Theorem 1.7.
+Section 3 introduces the local top-layer projection, defines the parity functional on Kameko’s $M_{1}$–summand, and proves Theorem 1.4.
+Finally, Section 4 discusses the extension to arbitrary algebraically closed base fields of characteristic 0, proves the base-change invariance of the local parity criterion, and combines it with the relevant arithmetic input to establish Theorem 1.6.
 2. Preliminaries
@@
 where $A^{>0,*}\subset A^{*,*}$ is the subspace of elements of positive total degree. The problem of describing a basis for the $QM_{n}^{d,*}(\mathbb{C})$ is the motivic hit problem.
-The quotient $N_{n}(\mathbb{C})$ is an $A^{\sharp}$–module, and there is an isomorphism
+Lemma 2.3 (Passage modulo $\tau$ preserves hit elements).
+Let
+$q_{\tau}:M_{n}(K)\longrightarrow N_{n}(K)=M_{n}(K)/(\tau)$
+be the quotient map. If
+$u\in A^{>0,*}\cdot M_{n}(K)\cap M_{n}(K)_{d,*},$
+then
+$q_{\tau}(u)\in A^{\sharp}_{+}(N_{n}(K))\cap N_{n}(K)_{d,*}.$
+Consequently, $q_{\tau}$ induces a surjective linear map
+$QM_{n}^{d,*}(K)\longrightarrow QN_{n}^{d,*}(K).$
+Proof.
+Write
+$u=\sum_{i}a_{i}\cdot v_{i}\qquad(a_{i}\in A^{>0,*},\ v_{i}\in M_{n}(K)).$
+Passing to the quotient modulo $\tau$, we obtain
+$q_{\tau}(u)=\sum_{i}\bar{a}_{i}\cdot q_{\tau}(v_{i}),$
+where $\bar{a}_{i}$ denotes the image of $a_{i}$ in
+$A^{\sharp}=A^{*,*}/(\tau).$
+If $\bar{a}_{i}=0$, the corresponding term vanishes. If $\bar{a}_{i}\neq 0$, then $\bar{a}_{i}$ still has positive total degree, hence $\bar{a}_{i}\in A^{\sharp}_{+}$. Therefore
+$q_{\tau}(u)\in A^{\sharp}_{+}(N_{n}(K)).$
+This proves the first assertion. The induced surjection on hit quotients follows immediately from the definition. ∎
+The quotient $N_{n}(\mathbb{C})$ is an $A^{\sharp}$–module. By Lemma 2.3, the quotient map modulo $\tau$ induces a surjective map on hit quotients; in fact, one has an isomorphism
 $QN_{n}^{d,*}(\mathbb{C}):=N_{n}(\mathbb{C})_{d,*}/\bigl(A^{\sharp}_{+}(N_{n}(\mathbb{C}))\cap N_{n}(\mathbb{C})_{d,*}\bigr)\cong QM_{n}^{d,*}(\mathbb{C})/\tau\,QM_{n}^{d,*}(\mathbb{C}),$
@@
 For any integers $n\geq 1$ and $1\leq k<n$, the monomial $z_{k}$ defined in (2.2) is not in $A^{>0,*}\cdot M_{n}(\mathbb{C})$.
-The key idea is to analyze the image of $Q_{0}$ in the quotient $N_{n}(\mathbb{C})/H_{n}$ for a certain subspace $H_{n}\subset N_{n}(\mathbb{C})$, and to describe the top layer of the hit quotient in terms of a pair of sets of monomials $M_{0}$ and $M_{1}$ which we review next.
+The key idea is to analyze the image of $Q_{0}$ in the quotient $N_{n}(\mathbb{C})/H_{n}$ for a certain subspace $H_{n}\subset N_{n}(\mathbb{C})$, and to describe the top layer in terms of two sets of monomials $M_{0}$ and $M_{1}$ which we review next.
 2.4. Weight vectors and the subspaces $F_{n}$, $G_{n}$, and $H_{n}$
@@
 $(\omega_{0}(z),\omega_{1}(z),\dots,\omega_{k}(z))<(k,n-1,\dots,n-1)$
-in the lexicographic order.
-He shows that the quotient $N_{n}/H_{n}$ decomposes as a direct sum over bigraded components $\Lambda^{a}_{n}\otimes(Y_{n}/G_{n})^{2b}$ with $a+2b=d$ and an explicit description of the contribution of $Q_{0}$ [3, Prop.
-4.1].
+in the lexicographic order. He shows that the quotient $(N_{n}/H_{n})_{d,*}$ decomposes as a direct sum of pieces
+$\Lambda_{n}^{a}\otimes(Y_{n}/G_{n})^{2b}/\pi\bigl(Q_{0}(\Lambda_{n}^{a+1}\otimes Y_{n}^{2(b-1)})\bigr)\qquad(a+2b=d),$
+see [3, Sec.
+4].
 We will only need the degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$.
-In this case, Kameko constructs two sets of monomials $M_{0}$ and $M_{1}$ in $\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ and shows that their images in $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis.
+In this case, Kameko constructs two sets of monomials $M_{0}$ and $M_{1}$ in $\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$ and shows that their images in $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis.
 Moreover, he proves that
@@
 We now turn to a more detailed analysis of this top summand.
-3. A parity functional on Kameko’s top layer
-In this section we work in the situation of the previous section with degree $d=k+2d_{1}$, $d_{1}=(n-1)(2^{k}-1)$, and monomial $z_{k}$ as in (2.2).
-We define a linear functional on the $\mathbb{F}_{2}$–span of the monomials in $M_{1}$ and show that its kernel exactly describes the intersection with the image of the motivic Steenrod algebra.
-3.1. The $M_{1}$–summand and its parity functional
-Fix integers $n\geq 2$ and $1\leq k<n$. Let $d_{1}=(n-1)(2^{k}-1)$ and $d=k+2d_{1}$. Following Kameko, we consider the summand
-$\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}\subset(N_{n}/H_{n})_{d,*},$
-and the sets $M_{0},M_{1}\subset\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ whose images form a basis of this summand. By [3, Eq. (4.3)], we have
-(3.1) $M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\},$
-and the images of $M_{0}\cup M_{1}$ in $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis.
-Let $\pi:Y_{n}\to Y_{n}/G_{n}$ be the quotient map, and extend it to
-$\pi:\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}\to\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
+3. A parity functional on Kameko’s local top layer
+In this section we work in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$.
+Our goal is to define a projection onto the $M_{1}$–summand before quotienting by hit elements and then to identify the image of hit elements under that projection.
+3.1. The local top-layer map and the parity functional
+Fix integers $n\geq 2$ and $1\leq k<n$. Let $d_{1}=(n-1)(2^{k}-1)$ and $d=k+2d_{1}$. Since
+$N_{n}^{d,*}=\bigoplus_{a+2b=d}\Lambda_{n}^{a}\otimes Y_{n}^{2b},$
+there is a canonical projection
+$\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$
+onto the summand with exactly $k$ exterior generators.
+Let
+$\pi:\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
+be the quotient map induced by $Y_{n}\to Y_{n}/G_{n}$.
+By Kameko’s description recalled above, there are subsets $M_{0},M_{1}\subset\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$ such that $\pi(M_{0})\cup\pi(M_{1})$ is a basis of $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ and
+(3.1) $M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\}.$
 Set
-$V:=\Bigl\langle\pi\bigl(\sigma(z_{k})\bigr)\,\Big|\,\sigma\in\mathrm{Mono}(k)\Bigr\rangle\subset\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
-Since the $\pi(\sigma(z_{k}))$ are part of a basis, they are linearly independent and form a basis of $V$.
-Definition 3.1 (Parity functional on the $M_{1}$–summand).
+$U_{0}:=\langle\pi(m)\mid m\in M_{0}\rangle,\qquad V:=\langle\pi(m)\mid m\in M_{1}\rangle.$
+Then
+$\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V.$
+Let
+$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
+be the projection with kernel $U_{0}$.
+Definition 3.1 (Local top-layer projection).
+Define the linear map
+$\vartheta:=p_{M_{1}}\circ\pi\circ\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow V.$
+We call $\vartheta(u)$ the local $M_{1}$–component (or local top-layer component) of $u\in N_{n}^{d,*}$.
+The point is that $\vartheta$ is defined directly on $N_{n}^{d,*}$.
+It records the $M_{1}$–part of an element after discarding only the lower-weight terms encoded by $G_{n}$ and the complementary basis block $M_{0}$; it does not mod out by hit elements.
+Definition 3.2 (Parity functional on the local top layer).
 List $\mathrm{Mono}(k)$ as $\{\sigma_{1},\dots,\sigma_{N}\}$ and set $m_{i}=\sigma_{i}(z_{k})$, so $M_{1}=\{m_{1},\dots,m_{N}\}$ with $N=\binom{n}{k}$.
@@
 $\varepsilon:V\to\mathbb{F}_{2},\qquad\varepsilon(v)=\sum_{i=1}^{N}c_{i}\pmod{2}$.
-Thus $\varepsilon(v)$ simply records whether the number of basis vectors $\pi(m_{i})$ that appear with coefficient $1$ in $v$ is even or odd.
+Thus $\varepsilon(v)$ records whether the number of basis vectors $\pi(m_{i})$ appearing in $v$ is even or odd.
 3.2. A linear algebra lemma
-We give a simple linear algebra fact that will be used repeatedly.
 Lemma 3.3.
@@
 We conclude that $E$ is an N1N-1–dimensional subspace of the even-parity hyperplane and coincides with it. ∎
-3.3. The $M_{1}$–part of the $Q_{0}$–image
-We now translate Kameko’s description of the image of $Q_{0}$ into a parity statement. Let $N_{n}=N_{n}(\mathbb{C})$ and $A^{\sharp}=A^{*,*}/(\tau)$ as before.
+3.3. The local $Q_{0}$–image in the $M_{1}$–summand
+We now translate Kameko’s description of the image of $Q_{0}$ into a parity statement for the local map $\vartheta$.
 Proposition 3.4 (Kameko [3, Props. 4.1, 5.1, 5.3]).
@@
   1. (1)
-    The quotient $(N_{n}/H_{n})_{d,*}$ decomposes as
-    $(N_{n}/H_{n})_{d,*}\cong\bigoplus_{a+2b=d}\Lambda^{a}_{n}\otimes(Y_{n}/G_{n})^{2b},$
-    and the summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ has basis given by the images of the monomials in $M_{0}\cup M_{1}$.
+    The set $\pi(M_{0}\cup M_{1})$ is a basis of
+    $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
+    In particular,
+    $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V.$
   2. (2)
-    For each monomial $z\in N_{n}$ of degree d1d-1 with $\omega_{0}(z)=k+1$, the image $\pi(Q_{0}(z))\in\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ is a linear combination of $\pi(z^{\prime})$ with $z^{\prime}\in M_{0}$ and elements of the form $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$ with $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$.
+    For each monomial $z\in N_{n}^{d-1,*}$ with $\omega_{0}(z)=k+1$, the element
+    $\pi\bigl(Q_{0}(z)\bigr)\in\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
+    is a linear combination of $\pi(z^{\prime})$ with $z^{\prime}\in M_{0}$ and elements of the form $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$ with $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$.
   3. (3)
-    There exist monotone permutations $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ (explicitly, the ones with $\sigma_{1}(1)=1,\sigma_{1}(2)=3,\dots,\sigma_{1}(k)=k+1$ and $\sigma_{2}(1)=2,\dots,\sigma_{2}(k)=k+1$) and a monomial $z\in N_{n}^{d-1,*}$ with $(\omega_{0}(z),\omega_{1}(z))=(k+1,n-2)$ and $\alpha_{i}(z)=k$ for all $i$, such that the $M_{1}$–part of $\pi(Q_{0}(z))$ is precisely $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$.
-The first two items are exactly Propositions 4.1 and 5.1 in [3]; the third is a reformulation of the second case in [3, Prop. 5.3] in our notation.
-We now apply Lemma 3.2.
-Proposition 3.4 (Parity of the $Q_{0}$–image in $V$).
-With notation as above, identify $V$ with $W=\mathbb{F}_{2}^{N}$ by mapping the basis $\pi(\sigma_{i}(z_{k}))$ to $e_{i}$. Then the $M_{1}$–part of the image of
-$Q_{0}:N_{n}^{d-1,*}\to N_{n}^{d,*}$
-is exactly the even-parity hyperplane
-$E=\{w\in W\mid w_{1}+\cdots+w_{N}=0\}\subset W.$
-Equivalently,
-$\varepsilon\bigl(\pi(Q_{0}(N_{n}^{d-1,*}))\cap V\bigr)=0.$
+    There exist monotone injections $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ (explicitly, the ones with $\sigma_{1}(1)=1,\sigma_{1}(2)=3,\dots,\sigma_{1}(k)=k+1$ and $\sigma_{2}(1)=2,\dots,\sigma_{2}(k)=k+1$) and a monomial $z\in N_{n}^{d-1,*}$ with $(\omega_{0}(z),\omega_{1}(z))=(k+1,n-2)$ and $\alpha_{i}(z)=k$ for all $i$, such that
+    $p_{M_{1}}\pi\bigl(Q_{0}(z)\bigr)=\pi\bigl(\sigma_{1}(z_{k})+\sigma_{2}(z_{k})\bigr).$
+We can now identify the local $Q_{0}$–image in $V$.
+Proposition 3.5 (Parity description of the local $Q_{0}$–image).
+With notation as above, identify $V$ with $W=\mathbb{F}_{2}^{N}$ by sending the basis $\pi(\sigma_{i}(z_{k}))$ to $e_{i}$. Then
+$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)=\ker(\varepsilon)\subset V.$
+Equivalently, the local $M_{1}$–part of the $Q_{0}$–image is exactly the even-parity hyperplane.
 Proof.
-By Proposition 3.3(2), the $M_{1}$–component of $\pi(Q_{0}(z))$ is always a linear combination of pairwise sums $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$, so the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ is contained in the subspace $E\subset W$ spanned by all $e_{i}+e_{j}$.
-Conversely, by Proposition 3.3(3) there exist $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ and a monomial $z_{0}\in N_{n}^{d-1,*}$ such that the $M_{1}$–part of $\pi(Q_{0}(z_{0}))$ is precisely $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$.
-Let $i_{0},j_{0}$ be the indices corresponding to $\sigma_{1},\sigma_{2}$, so that this $M_{1}$–part is the vector $e_{i_{0}}+e_{j_{0}}$ in $W$.
-Now let $S_{n}$ act on $N_{n}$ by permuting the variables $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$.
-This action preserves the subspaces $H_{n}$ and $G_{n}$, and it commutes with $Q_{0}$ and with the projection $\pi$, so for any permutation $\tau\in S_{n}$ we have
-$\pi(Q_{0}(\tau\cdot z_{0}))=\tau\cdot\pi(Q_{0}(z_{0})).$
-In particular, the $M_{1}$–part of $\pi(Q_{0}(\tau\cdot z_{0}))$ is $\pi(\tau\sigma_{1}(z_{k})+\tau\sigma_{2}(z_{k}))$, which corresponds to the vector $e_{i}+e_{j}$ where $i,j$ index the basis elements $\tau\sigma_{1}(z_{k})$ and $\tau\sigma_{2}(z_{k})$ of $V$.
-As $\tau$ varies, the pairs of kk–subsets $\tau\sigma_{1}(\{1,\dots,k\})$ and $\tau\sigma_{2}(\{1,\dots,k\})$ run through all pairs of kk–subsets differing in exactly one element; that is, through all edges of the Johnson graph $J(n,k)$ on the vertex set $\mathrm{Mono}(k)$.
-It is well-known that the Johnson graph $J(n,k)$ is connected for $1\leq k\leq n-1,$ between any two basis elements $e_{i}$ and $e_{j}$ there is a path $i=v_{0},v_{1},\dots,v_{t}=j$ in $J(n,k)$, and in characteristic $2$ we have
-$e_{i}+e_{j}=(e_{v_{0}}+e_{v_{1}})+(e_{v_{1}}+e_{v_{2}})+\cdots+(e_{v_{t-1}}+e_{v_{t}}).$
-Each summand $e_{v_{\ell-1}}+e_{v_{\ell}}$ is realized as the $M_{1}$–part of some $\pi(Q_{0}(z))$ as above, so $e_{i}+e_{j}$ lies in the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$.
-Thus all vectors $e_{i}+e_{j}$ are contained in that image.
-We conclude that the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ contains the subspace $E$ spanned by all $e_{i}+e_{j}$. Combined with the previous inclusion, this shows that it is exactly $E$.
-By Lemma 3.2, $E$ is the even-parity hyperplane in $W$. In particular, $\varepsilon$ vanishes on $\pi(Q_{0}(N_{n}^{d-1,*}))\cap V$. ∎
-3.4. The $A^{>0,*}$–image and the odd-parity gap
-We next extend Proposition 3.4 from the image of $Q_{0}$ to the image of the full motivic Steenrod algebra in the summand $V$.
-Lemma 3.5 (Reduced powers do not contribute to $M_{1}$).
-In the summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ of $(N_{n}/H_{n})_{d,*}$, the $M_{1}$–component of the image of the reduced powers $\mathcal{P}^{a}$ ($a\geq 1$) is zero.
-Equivalently, modulo $H_{n}$ and modulo $\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$, the only contribution to $V$ from $A^{\sharp}_{+}(N_{n})$ comes from $Q_{0}$.
+If $z\in N_{n}^{d-1,*}$ has $\omega_{0}(z)\neq k+1$, then $Q_{0}(z)$ lies in a summand $\Lambda_{n}^{a}\otimes Y_{n}^{2b}$ with $a\neq k$, hence $\operatorname{pr}_{k}(Q_{0}(z))=0$ and therefore $\vartheta(Q_{0}(z))=0$.
+Thus only monomials with $\omega_{0}(z)=k+1$ contribute.
+For such $z$, Proposition 3.4(2) shows that $\pi(Q_{0}(z))$ is a linear combination of $\pi(M_{0})$ and pairwise sums $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$.
+After applying $p_{M_{1}}$, we conclude that $\vartheta(Q_{0}(z))$ is a linear combination of vectors of the form $e_{i}+e_{j}$.
+Hence
+$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)\subseteq E,$
+where $E\subset W$ is the span of all $e_{i}+e_{j}$.
+Conversely, Proposition 3.4(3) provides one explicit edge $e_{i_{0}}+e_{j_{0}}$ in the image.
+Let $S_{n}$ act on $N_{n}$ by permuting the variables $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$.
+This action commutes with $Q_{0}$, preserves the subspace $G_{n}\subset Y_{n}$, and permutes the basis elements $\pi(\sigma(z_{k}))$ of $V$.
+Moreover, $S_{n}$ preserves $M_{0}$, since the defining condition for membership in $M_{0}$ is that $\alpha_{i}(z)<k$ for some $i$, and this condition is invariant under permutations of the indices.
+Hence $U_{0}=\langle\pi(m)\mid m\in M_{0}\rangle$ is $S_{n}$–stable, so the projection
+$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
+with kernel $U_{0}$ is $S_{n}$–equivariant. Hence, for every permutation $\tau\in S_{n}$,
+$\vartheta\bigl(Q_{0}(\tau\cdot z)\bigr)=p_{M_{1}}\pi\operatorname{pr}_{k}\bigl(Q_{0}(\tau\cdot z)\bigr)=\tau\cdot p_{M_{1}}\pi\operatorname{pr}_{k}\bigl(Q_{0}(z)\bigr)=\tau\cdot\vartheta\bigl(Q_{0}(z)\bigr).$
+Therefore all edges of the Johnson graph $J(n,k)$ occur in the image, because the orbit of the explicit pair $(\sigma_{1},\sigma_{2})$ under $S_{n}$ consists exactly of pairs of kk–subsets differing in one element.
+Since the Johnson graph $J(n,k)$ is connected for $1\leq k\leq n-1$, any vector $e_{i}+e_{j}$ is a sum of edge vectors along a path from $i$ to $j$.
+Thus every pairwise sum $e_{i}+e_{j}$ lies in $\vartheta(Q_{0}(N_{n}^{d-1,*}))$, so
+$E\subseteq\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
+Combining both inclusions gives
+$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)=E.$
+Finally, by Lemma 3.3, $E=\ker(\varepsilon)$. ∎
+3.4. From $Q_{0}$ to the full hit image in the local top layer
+We now extend Proposition 3.5 from the Bockstein to the full positive-degree part of the motivic Steenrod algebra.
+Lemma 3.6 (Reduced-power words vanish under $\vartheta$).
+Let $\mathcal{W}$ be the set of finite words in the generators $Q_{0},\mathcal{P}^{1},\mathcal{P}^{2},\dots$ of $A^{\sharp}$.
+Then $A^{\sharp}_{+}$ is spanned by the images of elements of $\mathcal{W}$.
+If $w\in\mathcal{W}$ contains some reduced power $\mathcal{P}^{c}$ with $c\geq 1$, then
+$\vartheta(w\cdot u)=0$
+for every $u\in N_{n}$ such that $w\cdot u\in N_{n}^{d,*}$. Consequently,
+$\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
 Proof.
-By construction, $G_{n}\subset Y_{n}$ is spanned by $P^{+}(Y_{n})$ and monomials of smaller weight, where $P$ is the subalgebra of $A^{\sharp}$ generated by the reduced powers $\mathcal{P}^{a}$ [3, Sec.
-4].
-In particular, $P^{+}(Y_{n})\subset G_{n}$ by definition.
-For $a\geq 1$, the operation $\mathcal{P}^{a}$ has bidegree $(2a,a)$ and acts trivially on $\Lambda_{n}$, so for any monomial $x\otimes y\in\Lambda^{k}_{n}\otimes Y_{n}^{2(d_{1}-a)}\subset N_{n}^{d-2a,*}$ we have
-$\mathcal{P}^{a}(x\otimes y)=x\otimes\mathcal{P}^{a}(y)\in\Lambda^{k}_{n}\otimes\bigl(P^{+}(Y_{n})\cap Y_{n}^{2d_{1}}\bigr)\subset\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}.$
-Here we used that $2(d_{1}-a)+2a=2d_{1}$, so the image lands in topological degree $d$ and polynomial degree $2d_{1}$.
-The summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ of $(N_{n}/H_{n})_{d,*}$ is the image of $\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ under the quotient map $\pi:Y_{n}\to Y_{n}/G_{n}$.
-Since $\mathcal{P}^{a}(x\otimes y)\in\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$ for all $x\otimes y$, its image under $\pi$ is zero.
-Thus the $M_{1}$–component of the image of every $\mathcal{P}^{a}$ ($a\geq 1$) in this summand vanishes.
-Equivalently, modulo $H_{n}$ and $\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$, the only nontrivial image in $V$ of $A^{\sharp}_{+}(N_{n})$ comes from $Q_{0}$.
-∎
-Definition 3.6 (The $M_{1}$–projection to $V$).
-Let $q:N_{n}\to N_{n}/H_{n}$ be the quotient map. In degree $d=k+2d_{1}$, Kameko’s decomposition [3, Prop. 4.1] gives a direct sum splitting of $(N_{n}/H_{n})_{d,*}$, hence a canonical projection
-$\operatorname{pr}_{k,d_{1}}:(N_{n}/H_{n})_{d,*}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
-Let
-$U_{0}:=\bigl\langle\pi(m)\mid m\in M_{0}\bigr\rangle\subset\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
-Since $\pi(M_{0})\cup\pi(M_{1})$ is a basis of $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$, we have a direct sum decomposition
-$\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V,$
-where $V=\langle\pi(M_{1})\rangle$ as in Section 3. Let
-$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
-be the projection with kernel $U_{0}$ (equivalently, $p_{M_{1}}$ kills $\pi(M_{0})$ and is the identity on $\pi(M_{1})$).
-Finally, define the composite linear map
-$\theta:N_{n}^{d,*}\xrightarrow{\ q\ }(N_{n}/H_{n})_{d,*}\xrightarrow{\ \operatorname{pr}_{k,d_{1}}\ }\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\xrightarrow{\ p_{M_{1}}\ }V.$
-We call $\theta(u)$ the $M_{1}$–component (in $V$) of $u$.
-Combining Proposition 3.4 and Lemma 3.5, we obtain the following.
-Proposition 3.7 (Parity vanishes on the $A^{\sharp}_{+}$–image in the $M_{1}$–summand).
-Let $u\in N_{n}^{d,*}$ be hit, i.e. $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$. Then the $M_{1}$–component $\theta(u)\in V$ has even parity:
-$\varepsilon(\theta(u))=0.$
-Equivalently,
-$\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)\subseteq\ker(\varepsilon)\subset V.$
-Proof.
-By Lemma 3.5, in the summand $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ the reduced powers $\mathcal{P}^{a}$ ($a\geq 1$) contribute trivially to the $M_{1}$–part.
-Therefore, after applying the projection $\theta$, the only possible contribution from $A^{\sharp}_{+}(N_{n})$ comes from the Bockstein $Q_{0}$.
-By Proposition 3.4, the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ is contained in the even-parity hyperplane $\ker(\varepsilon)\subset V$.
-Since $\theta$ is exactly the operation of passing to $(N_{n}/H_{n})_{d,*}$, projecting to the $(k,d_{1})$–summand, and then extracting the $M_{1}$–component, we obtain $\varepsilon(\theta(u))=0$ for every hit element $u$ of degree $d$.
-∎
+Since $A^{\sharp}_{+}$ is spanned by the images of words in $Q_{0},\mathcal{P}^{1},\mathcal{P}^{2},\dots$, it is enough by linearity to consider an operator represented by a single word $w$.
+Assume that $w$ contains at least one reduced power.
+The reduced powers act trivially on the exterior factor $\Lambda_{n}$ and act on $Y_{n}$ through the reduced-power subalgebra $P\subset A^{\sharp}$.
+Because the word $w$ contains a reduced power, the induced operator on the $Y_{n}$–factor lies in $P^{+}$.
+Hence, for every $u\in N_{n}$, the $Y_{n}$–part of $w\cdot u$ belongs to
+$P^{+}(Y_{n})\subset G_{n}.$
+The Bockstein $Q_{0}$ acts only on the exterior factor and does not alter the $Y_{n}$–part. Therefore, after applying the projection
+$\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}},$
+the result lies in
+$\Lambda_{n}^{k}\otimes G_{n}^{2d_{1}},$
+which is annihilated by
+$\pi:\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
+Hence $\vartheta(w\cdot u)=0$.
+Finally, since $Q_{0}^{2}=0$, the only nonzero word containing no reduced power is the single letter $Q_{0}$. The final equality follows. ∎
 We can now state the main structural theorem of this section.
-Theorem 3.8 (Theorem 1.4, General odd-parity non-hit theorem).
+Theorem 3.7 (Theorem 1.4).
 Let $n\geq 2$ and $1\leq k<n$, and let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$.
-Let $M_{1}$ and $V$ be as in (3.1), and let $\varepsilon:V\to\mathbb{F}_{2}$ be the parity functional from Definition 3.1.
-Let $\theta:N_{n}^{d,*}\to V$ be the $M_{1}$–projection from Definition 3.6.
+Let $V$ be the local $M_{1}$–summand defined above, let $\vartheta:N_{n}^{d,*}\to V$ be the local top-layer projection from Definition 3.1, and let $\varepsilon:V\to\mathbb{F}_{2}$ be the parity functional from Definition 3.2.
 Then:
@@
     One has an equality of subspaces of $V$:
-    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
+    $\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
   2. (2)
-    Consequently, $V/\theta(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*})$ is 11–dimensional over $\mathbb{F}_{2}$, generated by the class of any odd-parity element of $V$.
+    If $u\in N_{n}^{d,*}$ satisfies $\varepsilon(\vartheta(u))=1$, then $u\notin A^{\sharp}_{+}(N_{n})$.
   3. (3)
-    In particular, if $u\in\langle M_{1}\rangle\subset N_{n}^{d,*}$ has odd parity in its $M_{1}$–coordinates, then $u\notin A^{\sharp}_{+}(N_{n})$, so $u$ represents a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$.
+    In particular, if
+    $u=\sum_{\sigma\in S}\sigma(z_{k})\in N_{n}^{d,*}$
+    for some non-empty subset $S\subset\mathrm{Mono}(k)$ and $|S|$ is odd, then $u\notin A^{\sharp}_{+}(N_{n})$.
 Proof.
-By Proposition 3.7, we have the inclusion
-$\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)\subseteq\ker(\varepsilon).$
-For the reverse inclusion, Proposition 3.4 shows that for any $i\neq j$ the vector $e_{i}+e_{j}$ (i.e.
-any pairwise sum $\pi(\sigma_{i}(z_{k})+\sigma_{j}(z_{k}))$) occurs as the $M_{1}$–part of $\pi(Q_{0}(z))$ for some $z\in N_{n}^{d-1,*}$, hence lies in $\theta(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*})$.
-These vectors span $\ker(\varepsilon)$ by Lemma 3.2.
-Therefore,
-$\ker(\varepsilon)\subseteq\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr),$
-which proves (1).
-Since $\ker(\varepsilon)$ has codimension $1$ in $V$, statement (2) follows.
-Finally, if $u\in\langle M_{1}\rangle$ has odd parity, then $\theta(u)=\pi(u)\in V$ also has odd parity (because $\theta$ restricts to the identity on $\pi(M_{1})$ by construction).
-If $u$ were hit, then $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$ would imply $\theta(u)\in\ker(\varepsilon)$ by (1), contradicting odd parity.
-Hence $u$ is not hit and defines a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$, and therefore in $QM_{n}^{d,*}(\mathbb{C})$.
-∎
+By Lemma 3.6,
+$\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
+By Proposition 3.5, the latter space is exactly $\ker(\varepsilon)$. This proves (1).
+For (2), if $u$ were hit, then $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$, and so part (1) would imply $\vartheta(u)\in\ker(\varepsilon)$, contradicting $\varepsilon(\vartheta(u))=1$.
+For (3), note that if $u=\sum_{\sigma\in S}\sigma(z_{k})$, then $u\in\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$, its image under $\pi$ has no $M_{0}$–part, and therefore
+$\vartheta(u)=\sum_{\sigma\in S}\pi(\sigma(z_{k})).$
+Its parity is exactly $|S|$ modulo $2$. Hence odd cardinality implies $\varepsilon(\vartheta(u))=1$, and the claim follows from (2). ∎
+Remark 3.8.
+Theorem 3.7 identifies the image of hit elements after applying the local projection $\vartheta$.
+It does not assert that every even-parity linear combination of the monomials $\sigma(z_{k})$ is itself hit.
+The statement is weaker and sharper: every even-parity vector occurs as the local $M_{1}$–component of some hit element, while odd-parity vectors cannot occur in that way.
 Corollary 3.9 (A large family of non-hit classes).
@@
 Let $\bar{u}_{S}$ be the image of $u_{S}$ in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$.
-  1. (1)
 If $|S|$ is odd, then $\bar{u}_{S}$ is non-zero.
-  2. (2)
-    If $|S|$ is even, then the $M_{1}$–part of $\bar{u}_{S}$ lies in the image of $A^{\sharp}_{+}$, i.e. $\varepsilon(\bar{u}_{S})=0$.
 In particular, each individual translate $\sigma(z_{k})$ represents a non-zero class, and any sum of an odd number of translates does as well.
 Proof.
-The $M_{1}$–component of $u_{S}$ is precisely $\sum_{\sigma\in S}\pi(\sigma(z_{k}))$, whose parity is $|S|$ modulo $2$.
-If $|S|$ is odd, this parity is $1$, so by Theorem 3.8 the class of $u_{S}$ is non-zero in the hit quotient.
-If $|S|$ is even, the parity is 0 and the $M_{1}$–component lies in $\ker(\varepsilon)$, which is the $M_{1}$–part of the $A^{\sharp}_{+}$–image.
-∎
-4. Arithmetic of the Kameko degrees
-In this section we study the function $\beta(d)$ for the degrees $d=(n-1)(2^{k+1}-2)+k$ which arise in Kameko’s construction.
-The starting point is Proposition 2.1, which relates $\beta(d)>n$ to the inequality $\alpha(d+n)>n$.
-4.1. Kameko’s original estimates
-We recall Kameko’s computation from [3, Remark 2.2 and Proof of Prop. 1.5].
-Proposition 4.1 (Kameko [3, Remark 2.2]).
-Let $d=(n-1)(2^{k+1}-2)+k$.
-  1. (1)
-    If k=n1k=n-1, then $d+n=(n-1)\cdot 2^{n}+1$ and $\alpha(d+n)=\alpha(n-1)+1\leq n$.
-  2. (2)
-    If k=n2k=n-2, then $d+n=(n-1)\cdot 2^{n-1}$ and $\alpha(d+n)=\alpha(n-1)\leq n$.
-Hence $\beta(d)\leq n$ for k=n1k=n-1 and k=n2k=n-2.
-Proposition 4.2 (Kameko [3, Prop. 1.5]).
-Let $d=(n-1)(2^{k+1}-2)+k$ with k=n3k=n-3. Then
-$d+n=(n-1)\cdot 2^{n-2}-1=(n-2)\cdot 2^{n-2}+2^{n-3}+\cdots+2^{0}.$
-If $\alpha(n-2)\geq 3$, then $\alpha(d+n)>n$, and hence $\beta(d)>n$.
-Thus Kameko’s original family of motivic Peterson counterexamples arises from the case k=n3k=n-3. We now show that there are further families with k=n4k=n-4.
-4.2. A new family with k=n4k=n-4
-We continue to set $d=(n-1)(2^{k+1}-2)+k$ and $d_{1}=(n-1)(2^{k}-1)$, so that $d=k+2d_{1}$.
-Lemma 4.3.
-Let c=nkc=n-k. Then $1\leq c\leq n-1$ and
-$d+n=(n-1)\cdot 2^{n-c+1}+(2-c).$
-In particular, when $c=4$ (i.e. k=n4k=n-4) one has
-$d+n=(n-1)\cdot 2^{n-3}-2.$
-Proof.
-By (2.3) and k=nck=n-c we have
-$d$ $=(n-1)\bigl(2^{k+1}-2\bigr)+k$
-$=(n-1)\bigl(2^{n-c+1}-2\bigr)+(n-c)$
-$=(n-1)\cdot 2^{n-c+1}-2(n-1)+n-c.$
-Adding $n$ gives
-$d+n=(n-1)\cdot 2^{n-c+1}-2(n-1)+2n-c=(n-1)\cdot 2^{n-c+1}+(2-c),$
-as claimed. For $c=4$ this simplifies to
-$d+n=(n-1)\cdot 2^{n-3}-2.$
-We now specialize to a particular infinite sequence of $n$.
-Theorem 4.4 (Theorem 1.6, New numerical family with $\beta(d)>n$).
-Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let $d=(n-1)(2^{k+1}-2)+k$ as above. Then $\beta(d)>n$.
-Proof.
-We use Proposition 2.1. Set c=nkc=n-k. Since k=n4k=n-4, we have $c=4$, so by Lemma 4.3 we obtain
-$d+n=(n-1)\cdot 2^{n-3}-2.$
-Now assume $n=2^{r}+1$ with $r\geq 1$, so $n-1=2^{r}$. Then
-$d+n=2^{r}\cdot 2^{n-3}-2=2^{r+n-3}-2.$
-Let L=r+n3L=r+n-3. The binary expansion of $2^{L}-1$ is a string of $L$ ones, so $2^{L}-2=(2^{L}-1)-1$ has binary expansion $11\cdots 10$ with L1L-1 ones. Thus
-$\alpha(d+n)=\alpha(2^{L}-2)=L-1=(r+n-3)-1=r+n-4.$
-Since $n=2^{r}+1$, we have
-$\alpha(d+n)=r+2^{r}-3.$
-We compare this with $n=2^{r}+1$:
-$\alpha(d+n)>n\quad\Longleftrightarrow\quad r+2^{r}-3>2^{r}+1\quad\Longleftrightarrow\quad r-3>1\quad\Longleftrightarrow\quad r>4.$
-By hypothesis $r\geq 5$, so indeed $\alpha(d+n)>n$. By Proposition 2.1, this is equivalent to $\beta(d)>n$. ∎
-Remark 4.5.
-For small values of $r$ one can verify the above calculation by direct computation.
-For example, when $r=5$ we have $n=33$, $k=29$ and $d+n=2^{35}-2$, whose binary expansion has $34$ ones, so $\alpha(d+n)=34>33=n$.
-For $r=6$ we obtain $n=65$, $k=61$ and $d+n=2^{68}-2$ with $\alpha(d+n)=67>65$, and so on.
-Theorem 4.4 gives an infinite family of pairs $(n,k)$ with k=n4k=n-4 for which $\beta(d)>n$.
-In the next section we combine this with the odd-parity non-hit theorem to produce new motivic Peterson counterexamples.
-5. Base change and motivic counterexamples over general fields
+This is Theorem 3.7(3), interpreted in the hit quotient. ∎
+4. Base change and motivic Peterson counterexamples over general fields
 So far we have worked with the base field $\mathbb{C}$, as in Kameko’s original paper.
@@
 4.2. Extension of Kameko’s non-hit classes
-We now show that Kameko’s non-hit monomials and our parity refinements extend to arbitrary algebraically closed fields of characteristic 0.
+We now show that Kameko’s non-hit monomials and our local parity criterion extend to arbitrary algebraically closed fields of characteristic 0.
 Proposition 4.2.
@@
 Applying $\Phi_{K}$, we obtain $\Phi_{K}(z_{k})\in A^{>0,*}\cdot M_{n}(\mathbb{C})$.
-But by Proposition 2.3, $\Phi(z_{k})$ is not in this image.
+But by Proposition 2.4, $\Phi_{K}(z_{k})$ is not in this image.
 This is a contradiction.
@@
 ∎
-The same argument applies to any linear combination of the translates $\sigma(z_{k})$.
+The same argument applies to odd-parity linear combinations of the translates $\sigma(z_{k})$.
 Proposition 4.3.
@@
 $\overline{\Phi}_{K}:QN_{n}^{d,*}(K)\xrightarrow{\ \cong\ }QN_{n}^{d,*}(\mathbb{C}).$
-Under these identifications, the definitions of the subspaces $H_{n}$, $G_{n}$ and the sets $M_{0},M_{1}$ (hence also the subspace $V$ and the parity functional $\varepsilon$) are preserved.
-Moreover, the $M_{1}$–projection $\theta$ from Definition 3.6 is functorial with respect to $\Phi_{K}^{\sharp}$ (it is defined using the same direct-sum splitting and the same basis $\pi(M_{0})\cup\pi(M_{1})$), hence $\overline{\Phi}_{K}$ carries odd-parity classes to odd-parity classes.
-Therefore, if $u=\sum_{\sigma\in S}\sigma(z_{k})$ has odd parity, then its image $\overline{\Phi}_{K}(\bar{u})\in QN_{n}^{d,*}(\mathbb{C})$ is an odd-parity class, which is non-zero by Theorem 3.8 (over $\mathbb{C}$).
-Since $\overline{\Phi}_{K}$ is an isomorphism, $\bar{u}\neq 0$ in $QN_{n}^{d,*}(K)$.
+Under these identifications, the subspace $G_{n}$, the sets $M_{0},M_{1}$, the local top-layer summand $V$, the projection $\vartheta$, and the parity functional $\varepsilon$ are preserved, because they are defined by the same monomial formulas and the same weight conditions.
+Hence odd-parity classes over $K$ correspond exactly to odd-parity classes over $\mathbb{C}$.
+Therefore, if $u=\sum_{\sigma\in S}\sigma(z_{k})$ has odd parity, then its image in $QN_{n}^{d,*}(\mathbb{C})$ is non-zero by Theorem 3.7.
+Since $\overline{\Phi}_{K}$ is an isomorphism, the class of $u$ is already non-zero in $QN_{n}^{d,*}(K)$.
 Finally, using $QN_{n}^{d,*}(K)\cong QM_{n}^{d,*}(K)/\tau\,QM_{n}^{d,*}(K)$, we conclude that $u$ represents a non-zero class in $QM_{n}^{d,*}(K)$ as well.
@@
 4.3. Motivic Peterson counterexamples over $K$
-We now combine the arithmetic from Section 4 with the parity statement over general base fields.
+We now combine Kameko’s arithmetic input in the case k=n3k=n-3 and a direct binary calculation in the case k=n4k=n-4 with the local parity statement over general base fields.
 Theorem 4.4 (Motivic Peterson counterexamples for k=n3k=n-3 over $K$).
@@
 Proof.
-The inequality $\beta(d)>n$ is purely numerical and independent of the base field; it is proved in Proposition 4.2.
-For the non-vanishing of $QM_{n}^{d,*}(K)$, we may argue as follows.
-Over $\mathbb{C}$, Theorem 1.3 shows that $z_{k}$ is non-hit and has degree $d$.
-Hence $QM_{n}^{d,*}(\mathbb{C})\neq 0$.
-The isomorphism $M_{n}(K)\cong M_{n}(\mathbb{C})$ of $A^{*,*}$–modules induces an isomorphism
-$QM_{n}^{d,*}(K)\cong QM_{n}^{d,*}(\mathbb{C}),$
-so $QM_{n}^{d,*}(K)\neq 0$ as well. Alternatively, one can invoke Proposition 5.2, which implies that $z_{k}$ represents a non-zero class in $QM_{n}^{d,*}(K)$. ∎
-We now state and prove the analogue for the new family of degrees from Theorem 4.4.
+Assertion (1) is Kameko’s Proposition 1.5 [3], which states that if k=n3k=n-3 and $\alpha(n-2)\geq 3$, then $\beta(d)>n$.
+For the non-vanishing of $QM_{n}^{d,*}(K)$, over $\mathbb{C}$ the monomial $z_{k}$ is non-hit by Theorem 1.3, hence $QM_{n}^{d,*}(\mathbb{C})\neq 0$.
+Since $M_{n}(K)\cong M_{n}(\mathbb{C})$ as $A^{*,*}$–modules, it follows that $QM_{n}^{d,*}(K)\neq 0$.
+Equivalently, one may invoke Proposition 4.2.
+∎
 Theorem 4.5 (New motivic Peterson counterexamples at distance $4$).
@@
     represents a non-zero class in $QM_{n}^{d,*}(K)$.
-Thus these degrees $(n,d)$ form an infinite family of counterexamples to the motivic Peterson-type conjecture over $K$, distinct from the family with k=n3k=n-3.
+Thus these degrees $(n,d)$ form an infinite family of counterexamples to the motivic Peterson-type conjecture over $K$. This family is distinct from the case k=n3k=n-3.
 Proof.
-The inequality $\beta(d)>n$ follows from Theorem 4.4 and Proposition 2.1; this is purely number-theoretic and independent of $K$.
-For the non-vanishing of $QM_{n}^{d,*}(K)$ and the explicit non-hit elements, we use Proposition 5.3.
-In the present situation, the degree $d$ is again of the form $k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, so the odd-parity non-hit theorem applies.
-In particular, any $u_{S}$ with $|S|$ odd has odd parity in the $M_{1}$–summand and hence defines a non-zero class in $QN_{n}^{d,*}(K)$ and therefore in $QM_{n}^{d,*}(K)$.
-In particular, taking $S=\{\mathrm{id}\}$ shows that $z_{k}$ itself is non-hit and yields $QM_{n}^{d,*}(K)\neq 0$. This completes the proof. ∎
-Combining Theorems 5.4 and 5.5 proves Theorem 1.7 from the introduction.
+We first prove $\beta(d)>n$. By Proposition 2.1, it is enough to show that $\alpha(d+n)>n$. Since k=n4k=n-4, we compute
+$d+n$ $=(n-1)(2^{k+1}-2)+k+n$
+$=(n-1)(2^{n-3}-2)+(n-4)+n$
+$=(n-1)2^{\,n-3}-2(n-1)+2n-4$
+$=(n-1)2^{\,n-3}-2.$
+Now $n=2^{r}+1$, so $n-1=2^{r}$, and therefore
+$d+n=2^{r}\cdot 2^{\,n-3}-2=2^{\,r+n-3}-2.$
+Hence $d+n$ has binary expansion consisting of (r+n4)(r+n-4) ones followed by one zero, and thus
+$\alpha(d+n)=r+n-4.$
+Since $r\geq 5$, we have r+n4>nr+n-4>n, so $\alpha(d+n)>n$. Therefore $\beta(d)>n$ by Proposition 2.1.
+For the non-vanishing and the explicit classes, note that
+$d=k+2d_{1},\qquad d_{1}=(n-1)(2^{k}-1),$
+so Theorem 3.7 applies in this degree. Hence any odd-parity sum
+$u_{S}=\sum_{\sigma\in S}\sigma(z_{k})$
+with $|S|$ odd represents a non-zero class in $QN_{n}^{d,*}(K)$ and therefore in $QM_{n}^{d,*}(K)$ by Proposition 4.3.
+Taking $S=\{\mathrm{id}\}$ gives, in particular, that $z_{k}$ itself is non-hit, so $QM_{n}^{d,*}(K)\neq 0$.
+∎
+Combining Theorems 4.4 and 4.5 proves Theorem 1.6 from the introduction.
 Remark 4.6.